### v1

# A Photon Cloud Induced from an Axion Cloud

Abstract
It is known that the axion-photon coupling can lead to quantum stimulated emission of photons and classic exponential amplification of electromagnetic (EM) fields at half the axion mass frequency, when the axion density or the coupling constant is sufficiently large. In this work, we studied the EM photon cloud induced from an axion cloud around a Kerr black hole in the first order of the coupling constant classically. In the presence of a static EM background (like Wald extended solution) valid in realistic astrophysical environment, we found that an EM photon cloud emerges, oscillating at the same frequency as the axion cloud and growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied. Notably, the first-order study indicates that this photon cloud exists with no threshold for the coupling constant. The evolution of the EM photon cloud with time and azimuthal angle is obtained analytically while the cross-sectional distribution is solved numerically. Intriguingly, the induced EM field exhibits significantly different symmetries in contrast to the background EM field, which may serve as an indication of the existence of both an axion cloud and axion-photon coupling. Furthermore, in the near horizon region, the induced EM field has some overlap with the photon region as expected.
I Introduction
Axions or axion-like particles were introduced to solve the strong CP problem arising in QCD as it was found a discrepancy between the observed and predicted magnitude of the neutron electric dipole moment in the strong sector of the standard model [1, 2, 3]. Also axions appear in supersymmetric theories, and theories with extra dimensions including string theory from the compactification of the extra dimensions [4, 5, 6]. In (3+1)-dimensional case, a type called string-model independent axion [7, 8] is dual to the field strength of the Kalb-Ramond, spin-one, antisymmetric tensor field, and plays the role of a totally antisymmetric torsion in the geometry.
Recently there are extensive experimental and observational efforts to find axionic imprints. Superradiant mechanisms can give us information about their existance as axions form dense clouds around spinning black holes [9]. The superrradiance mechanism is a very well known method of extracting energy from a rotating black hole [10, 11, 12]. It was shown by Penrose [13] that if a particle accretes into a Kerr black hole then energy can be extracted from the black hole. It was shown by Misner [14] that there is a constrain of the frequency of scatter waves off the black hole to be less than the rotational frequency of the black hole. Then if this constrained condition is satisfied it was shown by Teukolsky [15] that this energy extraction mechanism also works in the case of EM and gravitational waves, though the Kerr solution remains stable under massless scalar, EM and gravitational perturbations without reflection.
It is well known that an isolated black hole cannot have an EM field unless it is endowed with a net electric charge. Therefore, for an isolated black hole to have some EM effects there should be a mechanism for charging it. If a black hole is in an astrophysical environment with a galactic magnetic field, EMs produced by external sources may charge it. In [16] an axisymmetric black hole which was placed in an originally uniform magnetic field and no significant field enhancement effects were found. However, if the black hole is rotating, the rotational effects of the black hole, produce electric fields near the black hole. Extraction of energy and charge from a black hole was studied in [17]. The superradiance effects and the multimessenger signals of an EM field were studied in [18, 19].
It was shown in [20] that the presence of light scalars like axions can trigger superradiant instabilities around Kerr BHs and lead to the deposition of the black hole rotation energy in a cloud. Also it was shown in [21] that the presence of an axion field can lead to the reversal of rotation in Kerr black holes. It was also shown that instabilities are generated and a fraction of the cloud energy is transferred to EM blasts. The resulting large energy density of the cloud rotates around the black hole, and results to a monochromatic gravitational wave radiation that drain the cloud over parametrically longer times [22].
Although the axion-photon coupling is expected to be very weak, it was suggested that the quantum stimulated emission of photons can be generated at large enough axion number [23, 24], and conjectured that blasts of light could be emitted from black hole systems [25, 26]. Recently at classical level, it was shown that EM fields can be exponentially amplified in Kerr-axion cloud system, when the amplitude of the axion cloud multiplied by the coupling constant is larger than a critical value [20]. Even in the presence of plasma, the EM instability is still controlled by the axionic coupling [9]. In [27], the origin of the instability from the axion-photon coupling was analyzed in flat space through the Mathieu equation, and it was found that a background EM field can produce an axion profile around a Kerr black hole perturbatively. On the other side, from the European Horizon Telescope (EHT) observation of the polarimetric measurements of M87*, stringent constraints on the axion-photon coupling constant were given in the mass range of $\sim\left(10^{-21}-10^{-20}\right)$eV from the azimuthal distribution of electric vector position angle [28]. Furthermore, the deviation of photon geodesics due to oscillating metric perturbations induced from the superradiant clouds was predicted in [11].
In this work we studied the EM photon cloud induced from an axion cloud via the axion-photon coupling surrounding a Kerr black hole. In the first order of the coupling constant, we found that a background EM field is necessary to have a first-order induced EM structure, and if it is stationary then the frequency of the induced photon cloud will be the same as the frequency of the axion cloud. It means that it grows or decays in accordance with the axion cloud when the superradiant condition for the axion field is satisfied or not, while the oscillating frequency is also the same. It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part. It is surprising that even with very small couplings, no threshold is needed, there exists an oscillating photon cloud at the same frequency as the axion cloud and it will grow exponentially when the axion cloud is growing. Whereas in the previous studies, without a background EM field, a threshold for the amplitude of the axion cloud multiplied by the coupling constant and an initial EM pulse are both required, otherwise there is no instability and the initial EM fluctuation will decay exponentially. In contrast, for the dark photon-photon coupling case, a dark photon cloud can transfer its energy to form a rotating EM field directly [18].
The work is organized as follows. Firstly in Sec. II we study the basic field equations in the first-order of the coupling constant, then in Sec. III we show how an EM photon cloud is induced from the axion cloud in the presence of an EM background and discuss symmetries. Moreover, in Sec. IV we study the induced electric and magnetic fields respectively, and in Sec. V we analyze the symmetries of most quantities and confirm the parity violation. Finally in Sec. VI we conclude.
II First-order Analysis of the Axion-Photon Coupling
In this work, we consider the interaction between an axion cloud and the EM field around a Kerr black hole with the action
$S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa}+\mathcal{L}_{m}\right]~{},$ (1)
$\mathcal{L}_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\mu_{1}^{2}\phi^{2}-\frac{k_{a}}{2}\phi F_{\mu\nu}{}^{*}F^{\mu\nu}~{},$ (2)
[…]
By variation of the above action, we obtain the basic field equations
$\nabla^{\mu}\nabla_{\mu}\phi-\mu_{1}^{2}\phi-\frac{k_{a}}{2}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~{},$ (3)
$\nabla_{\nu}F^{\mu\nu}+2k_{a}\nabla_{\nu}\phi{}^{*}F^{\mu\nu}=0~{},$ (4)
$R_{\mu\nu}-\frac{1}{2}g_{\mu\nu}R=\kappa\left(g_{\mu\nu}\mathcal{L}_{m}-2\frac{\delta\mathcal{L}_{m}}{\delta g^{\mu\nu}}\right)\equiv\kappa T_{\mu\nu}~{},$ (5)
with
$\frac{\delta\mathcal{L}_{m}}{\delta g^{\mu\nu}}=-\frac{1}{2}F^{\sigma}{}_{\mu}F_{\sigma\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla_{\nu}\phi-2k_{a}\phi F^{\sigma}{}_{\nu}{}^{*}F_{\sigma\mu}~{}.$ (6)
In our study, the Kerr geometry is fixed as the background, since the backreactions of both axion cloud and the EMs on the metric are in the second-order of the fluctuations. We expand the scalar and EM field equations to the first-order of the coupling constant $k_{a}$
$\nabla^{\mu}\nabla_{\mu}\phi_{(0)}-\mu_{1}^{2}\phi_{(0)}+k_{a}\left(\nabla^{\mu}\nabla_{\mu}\phi_{(1)}-\mu_{1}^{2}\phi_{(1)}-\frac{1}{2}F_{\mu\nu}^{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~{},$ (7)
$\nabla_{\nu}F^{\mu\nu}_{(0)}+k_{a}\left(\nabla_{\nu}F^{\mu\nu}_{(1)}+2\nabla_{\nu}\phi_{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~{},$ (8)
which clearly shows that at zero-order the axion cloud and the EM field obey their own field equations respectively, and at first-order the existence of a background EM field can give a substructure on the axion cloud while the existence of both axion cloud and a background EM field can induce an extra EM field in order of $k_{a}$. The first-order backreaction on the axion cloud has been studied in [27]. In this work, we focus on the EM field induced from the axion cloud in the presence of a background EM field.
The Kerr metric in Boyer-Lindquist (BL) coordinates is
$ds^{2}=-\frac{1}{\Sigma}\left(\Delta_{r}-a^{2}\sin^{2}{\theta}\right)dt^{2}+\frac{\Sigma}{\Delta_{r}}dr^{2}+\Sigma d\theta^{2}+\frac{\Gamma}{\Sigma}\sin^{2}{\theta}d\varphi^{2}-\frac{4aMr}{\Sigma}\sin^{2}{\theta}dtd\varphi~{},$ (9)
where
$\Sigma=r^{2}+a^{2}\cos^{2}{\theta}~{},$ (10)
$\Delta_{r}=r^{2}-2Mr+a^{2}~{},$ (11)
$\Gamma=\left(a^{2}+r^{2}\right)^{2}-\Delta_{r}a^{2}\sin^{2}{\theta}~{}.$ (12)
For the background EM field we consider the extended Wald solution
$A_{\mu}^{(0)}=A_{\mu}^{\rm Wald}=\frac{B_{0}\sin^{2}{\theta}}{2\Sigma}\left(-2aMr,0,0,\Gamma\right)~{},\quad A^{\mu}_{(0)}=A^{\mu}_{\rm Wald}=\left(0,0,0,\frac{B_{0}}{2}\right)~{},$ (13)
which describes the equilibrium state of the black hole absorbing charges and discharging processes from the surroundings of astrophysical environments such as the accretion disk, plasma and the interstellar medium [16]. It is worth noting that this extended Wald solution is still an exact solution in Kerr geometry.
For the axion cloud part, the bound states in Kerr metric can be obtained analytically in approximation when the fine structure constant $\alpha_{1}\equiv GM\mu_{1}/c=M\mu_{1}\ll 1$ in separable form as
$\phi_{(0)}=e^{-i\omega t+im\varphi}S_{lm}(\theta)R_{lm}(r)~{}$ (14)
with $\omega=\omega_{R}+i\omega_{I}$ that [29]
$\omega_{R}$ $\simeq$ $\mu_{1}\left(1-\frac{\alpha_{1}^{2}}{2\left(n+l+1\right)^{2}}\right)~{},$ (15)
$\omega_{I}$ $\simeq$ $\frac{1}{\gamma_{nlm}M}\left(\frac{am}{M}-2\mu_{1}r_{+}\right)\alpha_{1}^{4l+5}$ (16)
where $\gamma_{nlm}$ is a constant, $n$ is the principal quantum number, and $l$ corresponds to the orbital angular momentum quantum number. The condition for the superradiance to occur $\mu_{1}<\frac{am}{2Mr_{+}}=m\Omega_{H}$ is equivalent to $\alpha_{1}<\frac{am}{2r_{+}}=\frac{ma}{2M\left(1+\sqrt{1-a^{2}/M^{2}}\right)}$.
As we can see in eq. (8), since the background EM field $F_{\mu\nu}^{(0)}$ is independent of time $t$ and azimuthal angle $\varphi$, therefore the induced first-order EM field in form $F^{\mu\nu}_{(1)}\sim e^{-i\omega_{1}t+im_{1}\varphi}$ must have same frequency as the axion field (14) that $\omega_{1}=\omega$ and $m_{1}=m$. It means that the induced EM field grows or decays in accordance with the axion cloud when the superradiant condition for the axion field is satisfied or not, while the oscillating frequency is also the same. It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part.
It is necessary to clarify the difference of the EM frequencies. Turning back to the original Klein-Gordon equation (3), an axion field evolving as $\phi\sim e^{-i\omega t+im\varphi}$ would lead to EM field as $F_{\mu\nu}\sim e^{-i\omega_{EM}t+im_{EM}\varphi}$ at half frequency of the axion field $\omega_{EM}=\omega/2$. While if we look at the EM field equation (4), we cannot get a direct relation between the axion frequency and the EM frequency.
In quantum stimulated case [23], the half frequency was explained as the axion decay channel is into two photons resulting the photon energy at $E=m_{a}/2$. Also, the numerical evolution results [20, 9] gave EM oscillation frequency at half the axion mass frequency for supercritical couplings. In [27], the origin of the instability was analyzed in flat space through the Mathieu equation, which predicts instabilities for $\omega_{EM}=\mu_{1}/2,\mu_{1},3\mu_{1}/2,...$. However in [9] the frequency peaks at $\mu_{1},2\mu_{1},...$ were absent, and for high axionic couplings in dense plasmas $\omega_{p}>\mu_{1}/2$ the frequency centers at $\omega_{\rm EM}=\mu_{1}$ instead of the usual $\omega_{\rm EM}=\mu_{1}/2$, suggesting that the axion decay mechanism into two photons is forbidden and changes to process $\phi+\phi\to\gamma+\gamma$. Our understanding is that the half frequency appearing in classically numerical studies is because the main energy transfer from the axion cloud to EM field happens through the Klein-Gordon equation (3), so that even in curve spacetime we can still have a simple relation observed directly from the equation as we have analyzed previously.
III A Photon Cloud Induced from an axion cloud
Next, we try to solve the first-order induced EM field $F^{\mu\nu}_{(1)}$ in the Newman-Penrose (NP) formalism, in which using the Kinnersley’s null tetrad [30]
$l^{\mu}=\left(\frac{r^{2}+a^{2}}{\Delta_{r}},1,0,\frac{a}{\Delta_{r}}\right)~{},~{}n^{\mu}=\frac{1}{2\Sigma}\left(r^{2}+a^{2},-\Delta_{r},0,a\right)~{},~{}m^{\mu}=\frac{1}{\sqrt{2}\left(r+ia\cos{\theta}\right)}\left(ia\sin{\theta},0,1,\frac{i}{\sin{\theta}}\right)~{},$ (17)
we can construct three NP components
$\varphi_{0}=F_{\mu\nu}l^{\mu}m^{\nu}~{},\quad\varphi_{1}=\frac{1}{2}F_{\mu\nu}\left(l^{\mu}n^{\nu}+\overline{m}^{\mu}m^{\nu}\right)~{},\quad\varphi_{2}=F_{\mu\nu}\overline{m}^{\mu}n^{\nu}~{},$ (18)
where $\varphi_{0}$ is the ingoing radiation with spin weight $s=1$ and $\varphi_{2}$ represents outgoing radiation with spin weight $s=-1$. Note that there are only two real degree of freedoms though the three components are complex. The EM tensor can be reconstructed as
$F_{\mu\nu}=2\varphi_{0}\overline{m}_{[\mu}n_{\nu]}+2\overline{\varphi}_{0}m_{[\mu}n_{\nu]}+2\left(\varphi_{1}+\overline{\varphi}_{1}\right)n_{[\mu}l_{\nu]}+2\left(\varphi_{1}-\overline{\varphi}_{1}\right)m_{[\mu}\overline{m}_{\nu]}+2\varphi_{2}l_{[\mu}m_{\nu]}+2\overline{\varphi}_{2}l_{[\mu}\overline{m}_{\nu]}~{}.$ (19)
We derived the perturbation equation for electromagnetic field with sources in NP formalism [15]
$\left(D-2\rho\right)\varphi_{1}-\left(\bar{\delta}+\pi-2\alpha\right)\varphi_{0}+\kappa\varphi_{2}=-\frac{1}{2}J_{k}~{},$ (20)
$\left(\delta-2\tau\right)\varphi_{1}-\left(\Delta+\mu-2\gamma\right)\varphi_{0}+\sigma\varphi_{2}=-\frac{1}{2}J_{m}~{},$ (21)
$\left(D-\rho+2\varepsilon\right)\varphi_{2}-\left(\bar{\delta}+2\pi\right)\varphi_{1}+\lambda\varphi_{0}=-\frac{1}{2}J_{\bar{m}}~{},$ (22)
$\left(\delta-\tau+2\beta\right)\varphi_{2}-\left(\Delta+2\mu\right)\varphi_{1}+\nu\varphi_{0}=-\frac{1}{2}J_{\bar{m}}~{},$ (23)
and obtained the decoupled equation for the NP scalar $\varphi_{1}$
$\left[\Delta D+\left(\bar{\mu}-\mu-\gamma-\bar{\gamma}\right)D-\bar{\delta}\delta+\left(\bar{\tau}+2\alpha\right)\delta\right]\frac{\varphi_{1}}{\rho^{2}}=\left[\Delta+\left(\bar{\mu}-\mu-\gamma-\bar{\gamma}\right)\right]\frac{-J_{l}}{2\rho^{2}}+\left[\bar{\delta}-\left(\bar{\tau}+2\alpha\right)\right]\frac{-J_{m}}{2\rho^{2}}~{},$ (24)
where $D\equiv l^{a}\nabla_{a},\Delta\equiv n^{a}\nabla_{a},\delta\equiv m^{a}\nabla_{a},\bar{\delta}\equiv\overline{m}^{a}\nabla_{a}$ are derivative operators and $\kappa,\rho,\sigma,\tau,\nu,\mu,\lambda,\pi,\varepsilon,\beta,\gamma,\alpha$ are the spin coefficients. With the Kerr metric in BL coordinates (9), it can be written as
$\left[\frac{\left(r^{2}+a^{2}\right)^{2}}{\Delta_{r}}-a^{2}\sin^{2}{\theta}\right]\frac{\partial^{2}\psi}{\partial t^{2}}+\frac{4Mar}{\Delta_{r}}\frac{\partial^{2}\psi}{\partial t\partial\varphi}+\left[\frac{a^{2}}{\Delta_{r}}-\frac{1}{\sin^{2}{\theta}}\right]\frac{\partial^{2}\psi}{\partial\varphi^{2}}-\frac{\partial}{\partial r}\left(\Delta_{r}\frac{\partial\psi}{\partial r}\right)-\frac{1}{\sin{\theta}}\frac{\partial}{\partial\theta}\left(\sin{\theta}\frac{\partial\psi}{\partial\theta}\right)~{}$
$+\frac{2M\rho^{2}}{\bar{\rho}}\psi=\frac{1}{\bar{\rho}}\left\{\left[\Delta+\left(\bar{\mu}-\mu-\gamma-\bar{\gamma}\right)\right]\frac{J_{l}}{\rho^{2}}+\left[\bar{\delta}-\left(\bar{\tau}+2\alpha\right)\right]\frac{J_{m}}{\rho^{2}}\right\}~{},$ (25)
where $\psi=\varphi_{1}/\rho$, $J_{l}=J_{\mu}l^{\mu}$ and $J_{m}=J_{\mu}m^{\mu}$. The other two decoupled equations for $\varphi_{0}$ and $\varphi_{2}$ with sources can be found in [31].
In fact, in the scenario we are considering, the axion cloud is the source for the induced EM field that
$\nabla_{\nu}F^{\mu\nu}_{(1)}=-2\nabla_{\nu}\phi_{(0)}{}^{*}F^{\mu\nu}_{(0)}\equiv J^{\mu}_{(1)}~{}.$ (26)
We can take a good approximation to the bound states of axion cloud around a rotating black hole [12]
$\phi_{(0)}=\alpha_{1}^{2}\phi_{0}\frac{r}{M}e^{-\frac{r}{2M}\alpha_{1}^{2}}\cos{\left(\varphi-\omega_{R}t\right)}\sin{\theta}~{},$ (27)
which is valid for $\alpha\lesssim 0.2$ even at large black hole spin, and $\omega_{R}\simeq\mu_{1}$ is the oscillating frequency.
It can be separated into two parts
$\phi_{1}=e^{-i\omega t+i\varphi}R(r)S(\theta)~{},\quad\phi_{2}=e^{i\omega t-i\varphi}R(r)S(\theta)~{},$ (28)
with $R(r)=\alpha_{1}^{2}\phi_{0}\frac{r}{2M}e^{-\frac{r}{2M}\alpha_{1}^{2}}$ and $S(\theta)=\sin{\theta}$.
Using $\phi_{1}$ as the source and the NP scalars as partly separable variables
$\varphi_{01}=e^{-i\omega_{1}t+im_{1}\varphi}f_{01}\left(r,\theta\right)~{},~{}\varphi_{11}=e^{-i\omega_{1}t+im_{1}\varphi}f_{11}\left(r,\theta\right)~{},~{}\varphi_{21}=e^{-i\omega_{1}t+im_{1}\varphi}f_{21}\left(r,\theta\right)~{},$ (29)
we can obtain the three decoupled differential equations respectively for $f_{01}\left(r,\theta\right)$, $f_{11}\left(r,\theta\right)$ and $f_{21}\left(r,\theta\right)$ after dropping off the time $t$ and azimuthal angle $\varphi$ dependence by $\omega_{1}=\omega$ and $m_{1}=m=1$. We can also do that for $\phi_{2}$.
We replace $\omega=\mu_{1}=\alpha_{1}/M$, $a_{0}\equiv a/M$, $r_{0}\equiv r/M$, and define new dimensionless quantities $f_{01M}\equiv Mf_{01}\left(r,\theta\right)$, $f_{11M}\equiv Mf_{11}\left(r,\theta\right)$, $f_{\rm 21M}\equiv Mf_{21}\left(r,\theta\right)$. After absorbing some coefficients from the source part into the coupling constant $k_{a}$, we now use a new dimensionless coupling constant $\varepsilon\equiv\alpha_{1}^{2}B_{0}M\phi_{0}k_{a}$ as the small quantity for expansion, then the new field equations are dimensionless
${\rm eq}f_{\rm 01M}:\left\{\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 01M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (30)
${\rm eq}f_{\rm 11M}:\left\{\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 11M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (31)
${\rm eq}f_{\rm 21M}:\left\{\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 21M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (32)
and there are only two dimensionless parameters $a_{0}$ and $\alpha_{1}$.
Different from the quasi-normal modes and bound states, here the frequency of the induced EM field is already fixed by the frequency of the source (axion cloud), it is not an eigenvalue problem and no boundary condition should be imposed. Actually we can obtain the asymptotic behaviors of the functions at the boundaries by expanding the equations (we choose decaying solutions at spatial infinity), and then solve the equations numerically.
With expansions and the regular condition at poles $\theta=0,\pi$, the asymptotic behaviors at the boundaries $r_{0}=r_{\rm h0}\equiv 1+\sqrt{1-a_{0}^{2}}$, $r_{0}\to\infty$, $\theta=0,\pi$ are as follows
$f_{\rm 01M}\left(r_{\rm h0},\theta\right)=\frac{2\sqrt{2}a_{0}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{\theta}\sin{2\theta}}{\left(2\alpha_{1}r_{h0}-a_{0}-2i\sqrt{1-a_{0}^{2}}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (33)
$f_{\rm 11M}\left(r_{\rm h0},\theta\right)=\frac{2i\sqrt{1-a_{0}^{2}}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{2\theta}}{\left(a_{0}-2\alpha_{1}r_{h0}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (34)
$f_{\rm 21M}\left(r_{\rm h0},\theta\right)=0~{},$ (35)
$f_{\rm 01M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (36)
$f_{\rm 11M}\left(r_{0}\to\infty,\theta\right)=-\frac{(\alpha_{1}+3i)\sin{2\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{4\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (37)
$f_{\rm 21M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{2\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (38)
$f_{\rm 01M}\left(r_{0},0\right)=0,\quad f_{\rm 11M}\left(r_{0},0\right)=0,\quad f_{\rm 21M}\left(r_{0},0\right)=0,$ (39)
$f_{\rm 01M}\left(r_{0},\pi\right)=0,\quad f_{\rm 11M}\left(r_{0},\pi\right)=0,\quad f_{\rm 21M}\left(r_{0},\pi\right)=0~{},$ (40)
where the outgoing part $f_{\rm 21M}\left(r_{0},\theta\right)$ just vanishes at the horizon in accord with the fact that there is no outgoing radiation at the event horizon. It should also result from the fact that $\varphi_{2}$ is with negative spin weight which can not retrograde against the rotation at event horizon.
In addition, it is clear that the asymptotic behaviors for $f_{\rm 01M}$ and $f_{\rm 11M}$ at event horizon $r_{\rm h0}$ are only valid when $\alpha_{1}\neq\frac{a_{0}}{2r_{\rm h0}}$, which is exactly the critical condition for whether superradiance occurs. In this critical case with $\alpha_{1}=\frac{a_{0}}{2r_{\rm h0}}$, the imaginary part $\omega_{I}$ disappears in the original field equation for the axion field, so that the whole frequency is not complex and the original boundary condition at event horizon is not valid, in which case another regular condition at origin should be applied. Here we do not go further on this case.
Similarly, for the NP scalars induced from $\phi_{2}$
$\varphi_{02}=e^{i\omega_{1}t-im_{1}\varphi}f_{02}\left(r,\theta\right)~{},~{}\varphi_{12}=e^{i\omega_{1}t-im_{1}\varphi}f_{12}\left(r,\theta\right)~{},~{}\varphi_{22}=e^{i\omega_{1}t-im_{1}\varphi}f_{22}\left(r,\theta\right)~{},$ (41)
we can get the corresponding equations and boundary conditions for $f_{02M}\left(r_{0},\theta\right)$, $f_{12M}\left(r_{0},\theta\right)$, $f_{22M}\left(r_{0},\theta\right)$.
From the analysis of the equations, we found that the induced NP scalars from the two parts $\phi_{1}$ and $\phi_{2}$ of the axion cloud have following relations (here we remove the coefficient $M$)
${\rm Re}\left[f_{01}\left(r,\theta\right)\right]={\rm Re}\left[f_{02}\left(r,\pi-\theta\right)\right]~{},\quad{\rm Im}\left[f_{01}\left(r,\theta\right)\right]=-{\rm Im}\left[f_{02}\left(r,\pi-\theta\right)\right]~{},$ (42)
${\rm Re}\left[f_{11}\left(r,\theta\right)\right]=-{\rm Re}\left[f_{12}\left(r,\pi-\theta\right)\right]~{},\quad{\rm Im}\left[f_{11}\left(r,\theta\right)\right]={\rm Im}\left[f_{12}\left(r,\pi-\theta\right)\right]~{},$ (43)
${\rm Re}\left[f_{21}\left(r,\theta\right)\right]={\rm Re}\left[f_{22}\left(r,\pi-\theta\right)\right]~{},\quad{\rm Im}\left[f_{21}\left(r,\theta\right)\right]=-{\rm Im}\left[f_{22}\left(r,\pi-\theta\right)\right]~{},$ (44)
which means that with the above transformations, we can start from one set of the equations to the other set. These symmetries from analysis are consistent with the numerical results, thus here we only exhibit the distribution for those from $\phi_{1}$ in FIG. 1.
Figure 1: The real and imaginal parts of $f_{01M}\left(r_{0},\theta\right)$, $f_{11M}\left(r_{0},\theta\right)$, $f_{21M}\left(r_{0},\theta\right)$ induced from $\phi_{1}$ are plotted with $a_{0}=0.5$ and $\alpha_{1}=0.1$ in the scale of $200M$, where the superradiant condition is satisfied $\alpha_{1}<\frac{a_{0}}{2r_{\rm h0}}$.
From FIG. 1 we can see that, the ingoing radiation $f_{01M}$ is slightly larger than the outgoing radiation $f_{21M}$, and the residential part $f_{11M}$ presents a ripple-like pattern. Usually the residential $\varphi_{1}$ part is neglected, since its equation is not separable even without any source and it is not related to ingoing or outgoing part. But here, it is shown that the imaginary part of $f_{11M}$ gets much larger amplitude than the other parts, indicating that the residential part is the primary constituent of the EM photon cloud we are studying. In fact, both $\varphi_{0}$ and $\varphi_{2}$ contain ingoing and outgoing parts, people assign ingoing/outgoing radiation to $\varphi_{0}$/$\varphi_{2}$ due to which one is dominant at spatial infinity.
Subsequently, we can construct the EM tensor induced from $\phi_{1}$ using (19)
$F_{\mu\nu}^{1}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]-4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]~{},$ (45)
$P_{\mu\nu}^{1}\left(r,\theta\right)\equiv f_{01}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{11}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{11}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{21}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{}.$ (46)
Similarly, for the EM tensor induced from $\phi_{2}$ we have
$F_{\mu\nu}^{2}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]+4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]~{},$ (47)
$P_{\mu\nu}^{2}\left(r,\theta\right)\equiv f_{02}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{12}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{12}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{22}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{},$ (48)
then the whole EM field induced from the axion cloud $\phi_{(0)}=\phi_{1}+\phi_{2}$ evolves as
$F_{\mu\nu}^{(1)}=F_{\mu\nu}^{1}+F_{\mu\nu}^{2}=\cos{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}\left(r,\theta\right)~{},$ (49)
$F_{\mu\nu}^{\rm cos}\left(r,\theta\right)=4~{}{\rm Re}\left[\left(f_{01}+f_{02}\right)\overline{m}_{[\mu}n_{\nu]}+\left(f_{11}+f_{12}\right)\left(n_{[\mu}l_{\nu]}+m_{[\mu}\overline{m}_{\nu]}\right)+\left(f_{21}+f_{22}\right)l_{[\mu}m_{\nu]}\right]~{},$ (50)
$F_{\mu\nu}^{\rm sin}\left(r,\theta\right)=4~{}{\rm Im}\left[\left(f_{02}-f_{01}\right)\overline{m}_{[\mu}n_{\nu]}+\left(f_{12}-f_{11}\right)\left(n_{[\mu}l_{\nu]}+m_{[\mu}\overline{m}_{\nu]}\right)+\left(f_{22}-f_{21}\right)l_{[\mu}m_{\nu]}\right]~{}.$ (51)
So far we have considered an EM photon cloud sourced by an axion cloud surrounding a Kerr black hole which is described by (49). This is one of the main results of our work. Moreover, it means that for an arbitrary current $J^{a}=e^{\pm i\omega t\mp im\varphi}S\left(r,\theta\right)$ as the source we will always get a photon cloud evolving as
$F_{\mu\nu}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}\left(r,\theta\right)\right]\pm 4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}\left(r,\theta\right)\right]$.
In fact, even for photons without any couplings, they can have an effective mass given by plasma and make superradiance to happen [32, 33].
IV The induced electric and magnetic fields
In this Section we will study the induced electric field and magnetic field respectively. Considering the $(3+1)$-decomposition of the spacetime [34]
$ds^{2}=-\alpha^{2}dt^{2}+\gamma_{ij}\left(dx^{i}+\beta^{i}dt\right)\left(dx^{j}+\beta^{j}dt\right)~{},$ (52)
where $\beta^{i}$ is shift vector, $\gamma_{ij}$ is the spatial metric and the lapse function $\alpha=\sqrt{\frac{\Delta_{r}\Sigma}{\Delta_{r}\Sigma+2Mr\left(a^{2}+r^{2}\right)}}$ in Kerr metric (9). We can define the electric field and magnetic field [35]
$E^{a}:=n_{b}F^{ab}~{},\quad B^{a}:=n_{b}{}^{*}F^{ab}~{},$ (53)
with respect to an Eulerian observer $n^{a}=\left(1/\alpha,-\beta^{i}/\alpha\right)$ whose worldline is orthogonal to the spacelike hypersurface.
Therefore we have
$E^{i}_{(1)}=\cos{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}E^{i}_{\rm sin}\left(r,\theta\right)~{},$ (54)
$B^{i}_{(1)}=\cos{\left(m\varphi-\omega t\right)}B^{i}_{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}B^{i}_{\rm sin}\left(r,\theta\right)~{},$ (55)
where
$E^{i}_{\rm cos}=-\alpha F^{i0}_{\rm cos}~{},\quad E^{i}_{\rm sin}=-\alpha F^{i0}_{\rm sin}~{},$ (56)
$B^{i}_{\rm cos}=-\frac{1}{2}\alpha\epsilon^{i0jk}F_{jk}^{\rm cos}~{},\quad B^{i}_{\rm sin}=-\frac{1}{2}\alpha\epsilon^{i0jk}F_{jk}^{\rm sin}~{}.$ (57)
In FIG. 2-4, we plot the components of induced electric and magnetic fields in cosine and sine sections respectively, in contrast with the background Wald case. In some panels with large scale like $100M$, we use the dotted circles to denote the radius $r_{\rm max}/M=2/\alpha_{1}^{2}$ of the maximum value of axion cloud, while in some panels with small scale like $5M$, we compare with the boundary of photon region (dotdashed curves) and event horizon (dashed circles). It is a good case to compare the distribution of the EMs (in field perspective) with the photon region (spherical photon orbits in particle perspective).
Figure 2: The components of the induced electric fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
Figure 3: The components of the induced magnetic fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
From the figures we can see that, the radial components $E^{r}_{\rm cos}$ and $E^{r}_{\rm sin}$ of the induced electric field both exhibit quadrupole moments with a parity violation ($P=-1$). In contrast, the radial component of electric field of the background Wald case is in dipole moments with a parity symmetry ($P=+1$). While the $\theta$ components $E^{\theta}_{\rm cos}$ and $E^{\theta}_{\rm sin}$ of the induced electric field both exhibit a parity symmetry ($P=+1$). In contrast, the $\theta$ component of electric field of the background Wald case is in quadrupole moments with a parity violation ($P=-1$). For the $\varphi$ components $E^{\varphi}_{\rm cos}$ and $E^{\varphi}_{\rm sin}$ of the induced electric field, they both show a parity violation symmetry ($P=-1$). Conversely, the symmetries for the components of magnetic fields are all opposite with those for electric fields. Besides, we can observe some evident coincides between the distribution of the induced electric and magnetic fields and the boundary of photon region.
Note that both electric and magnetic fields of the background Wald case do not contain $\varphi$ components, and they only distribute outside the ergoregion, while the induced EM field can live within the ergoregion. This can be understood as “no static observer is allowed inside the ergoregion”.
As we know, superradiance only happens for massive fields, where the mass term makes them gathering around the black hole and sustainably dissipate near the event horizon. The dissipation is a crucial ingredient for superradiance, in fact even for a rotating compact object without a horizon, superradiance can occur via other dissipation mechanism, like friction [10].
While for the axion-photon coupling case, due to the antisymmetry of $\epsilon^{\mu\nu\rho\sigma}$ there is never an effective mass term for the photons, though sometimes people call it “tachyonic-like instability” [20]. Also it has been mentioned that the time-dependent axion is a crucial ingredient to trigger this instability when $k_{a}\phi_{0}$ is larger than a critical value.
In our first-order study with an EM background, even for small couplings, the exponential growing happens for the EM field coupled to the axion field, as we have discussed in Sec. II. From our perspective, the reason is that the stationary axion cloud ties the EM field onto the nearby of the black hole and serves as a source continuously, which makes the EM field dissipate sustainably from the axion cloud to the photon region and finally to the event horizon. As it is shown in the figures, the induced EM field does not concentrate around the maximum value of axion cloud but its main distribution is inside the dotted circles. In other words, the most dense places for the photon cloud do not relate to the largest oscillating position (same radius $r_{\rm max}$) of the axion cloud but inside it.
Figure 4: The components of the electric and magnetic fields are plotted for the background Wald case, where we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
V Analysis of symmetries and parity violation
It is well known that the term $F_{\mu\nu}{}^{*}F^{\mu\nu}$ violates parity and couples to the pseudo-scalar axion field, making the whole term $\phi F_{\mu\nu}{}^{*}F^{\mu\nu}$ parity invariant and leading to CP violation.
Here in curved spacetime and classically, we obtain the same symmetry for the parity violation ($P=-1$). From (49), we have
$F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}~{}$
$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}~{},$
where $m=1$ and
$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}$ $=$ $8\left[({\rm Im}[f_{01}]+{\rm Im}[f_{02}])({\rm Re}[f_{21}]+{\rm Re}[f_{22}])+({\rm Re}[f_{01}]+{\rm Re}[f_{02}])({\rm Im}[f_{21}]+{\rm Im}[f_{22}])\right.$ (59)
$\left.-2({\rm Im}[f_{11}]+{\rm Im}[f_{12}])({\rm Re}[f_{11}]+{\rm Re}[f_{12}])\right]~{},$
$F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ $=$ $-8({\rm Im}[f_{01}]-{\rm Im}[f_{02}])({\rm Re}[f_{21}]-{\rm Re}[f_{22}])-8({\rm Re}[f_{01}]-{\rm Re}[f_{02}])({\rm Im}[f_{21}]-{\rm Im}[f_{22}])$ (60)
$+16({\rm Im}[f_{11}]-{\rm Im}[f_{12}])({\rm Re}[f_{11}]-{\rm Re}[f_{12}])~{},$
$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[-{\rm Im}[f_{01}]{\rm Im}[f_{21}]+{\rm Re}[f_{01}]{\rm Re}[f_{21}]+{\rm Im}[f_{02}]{\rm Im}[f_{22}]-{\rm Re}[f_{02}]{\rm Re}[f_{22}]\right.$ (61)
$\left.+{\rm Im}[f_{11}]^{2}-{\rm Re}[f_{11}]^{2}-{\rm Im}[f_{12}]^{2}+{\rm Re}[f_{12}]^{2}\right]~{},$
all exhibit a sign change for $\theta\to\pi-\theta$, in agreement with the results in FIG. 5. Combined with no sign change for $\varphi\to\varphi+\pi$, we know that there is a parity violation $P=-1$ for $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$.
In addition, it shows that except for the evolving parts with a double frequency $2\omega$, there are also residential parts that are time-independent.
Figure 5: The components $M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ of the parity violation term $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
Similarly, the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ evolves as
$F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}~{}$
$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}~{},$
where $m=1$ and
$F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}$ $=$ $8\left[-({\rm Im}[f_{01}]+{\rm Im}[f_{02}])({\rm Im}[f_{21}]+{\rm Im}[f_{22}])+({\rm Re}[f_{01}]+{\rm Re}[f_{02}])({\rm Re}[f_{21}]+{\rm Re}[f_{22}])\right.~{}$ (63)
$\left.+({\rm Im}[f_{11}]+{\rm Im}[f_{12}])^{2}-({\rm Re}[f_{11}]+{\rm Re}[f_{12}])^{2}\right]~{},$
$F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[({\rm Im}[f_{01}]-{\rm Im}[f_{02}])({\rm Im}[f_{21}]-{\rm Im}[f_{22}])-({\rm Re}[f_{01}]-{\rm Re}[f_{02}])({\rm Re}[f_{21}]-{\rm Re}[f_{22}])\right.$ (64)
$\left.-({\rm Im}[f_{11}]-{\rm Im}[f_{12}])^{2}+({\rm Re}[f_{11}]-{\rm Re}[f_{12}])^{2}\right]~{},$
$F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[-{\rm Im}[f_{01}]{\rm Re}[f_{21}]-{\rm Re}[f_{01}]{\rm Im}[f_{21}]+{\rm Im}[f_{02}]{\rm Re}[f_{22}]+{\rm Re}[f_{02}]{\rm Im}[f_{22}]\right.$ (65)
$\left.+2{\rm Im}[f_{11}]{\rm Re}[f_{11}]-2{\rm Im}[f_{12}]{\rm Re}[f_{12}]\right]~{},$
all keep a parity symmetry $P=+1$ under $\theta\to\pi-\theta$, in agreement with the results in FIG. 6. Also there is no sign change under $\varphi\to\varphi+\pi$, hence the whole $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ keeps parity symmetry $P=+1$.
We know that the null conditions for EM fields are $F_{\mu\nu}F^{\mu\nu}=0$ and $F_{\mu\nu}{}^{*}F^{\mu\nu}=0$, which is equivalent to $\varphi_{0}\varphi_{2}-\varphi_{1}^{2}=0$. Here with the axion-photon coupling, it shows obviously that the null conditions are not satisfied anymore.
Figure 6: The components $M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ of the contraction of the EM field strength tensor $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
Moreover, we study the effects of the two parameters $a_{0},~{}\alpha_{1}$ and find that the spin parameter $a_{0}$ only has slight impacts on both the amplitude and the structure of the distribution for each components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$, while these quantities are sensitive to the fine structure constant $\alpha_{1}$. Therefore, in FIG. 5-6 we plot the figures for the components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ with fixed $a_{0}=0.8$ and different values of $\alpha_{1}$. From the figures we can see that, for all the components, the increase of $\alpha_{1}$ leads to a significant decrease in amplitude and a slight concentration on the distribution, which indicates that for fixed black hole mass $M$ the smaller axion mass $\mu_{1}$ (larger reduced Compton wavelength) makes the interaction with photons more efficient and stronger. Also, the slight concentration on the distribution should result from the concentration of the axion cloud with larger $\alpha_{1}$, where larger axion mass $\mu_{1}$ (smaller reduced Compton wavelength) makes the axion cloud concentrate. In general, the induced EM scalars $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ concentrate in the sine parts $M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ and $M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$.
For the electric and magnetic fields respectively, applying the relations (42)-(44), we can obtain some symmetries for the components in cosine and sine parts
$E^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{r}_{\rm cos}\left(r,\theta\right)~{},\quad B^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{r}_{\rm cos/sin}\left(r,\theta\right)~{},$ (66)
$E^{\theta}_{\rm cos/sin}\left(r,\pi-\theta\right)=E^{\theta}_{\rm cos/sin}\left(r,\theta\right)~{},\quad B^{\theta}_{\rm cos/sin}\left(r,\pi-\theta\right)=-B^{\theta}_{\rm cos/sin}\left(r,\theta\right)~{},$ (67)
$E^{\varphi}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{\varphi}_{\rm cos/sin}\left(r,\theta\right)~{},\quad B^{\varphi}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{\varphi}_{\rm cos/sin}\left(r,\theta\right)~{},$ (68)
which can be clearly seen in FIG. 2-3. We further have
$E^{r}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=E^{r}_{(1)}\left(t,r,\theta,\varphi\right)~{},\quad B^{r}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-B^{r}_{(1)}\left(t,r,\theta,\varphi\right)~{},$ (69)
$E^{\theta}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-E^{\theta}_{(1)}\left(t,r,\theta,\varphi\right)~{},\quad B^{\theta}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=B^{\theta}_{(1)}\left(t,r,\theta,\varphi\right)~{},$ (70)
$E^{\varphi}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=E^{\varphi}_{(1)}\left(t,r,\theta,\varphi\right)~{},\quad B^{\varphi}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-B^{\varphi}_{(1)}\left(t,r,\theta,\varphi\right)~{},$ (71)
and the background EM field transforms as
$E^{r}_{(0)}\left(t,r,\pi-\theta,\varphi+\pi\right)=E^{r}_{(0)}\left(t,r,\theta,\varphi\right)~{},\quad B^{r}_{(0)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-B^{r}_{(0)}\left(t,r,\theta,\varphi\right)~{},$ (72)
$E^{\theta}_{(0)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-E^{\theta}_{(0)}\left(t,r,\theta,\varphi\right)~{},\quad B^{\theta}_{(0)}\left(t,r,\pi-\theta,\varphi+\pi\right)=B^{\theta}_{(0)}\left(t,r,\theta,\varphi\right)~{}.$ (73)
These transformations are equivalent to $\overset{\rightharpoonup}{E}\rightarrow-\overset{\rightharpoonup}{E}$ and $\overset{\rightharpoonup}{B}\rightarrow\overset{\rightharpoonup}{B}$ under $\overset{\rightharpoonup}{r}\rightarrow-\overset{\rightharpoonup}{r}$. Therefore the induced electric field still keep a parity symmetry with $P=-1$ as a vector, and the induced magnetic field keeps the parity symmetry with $P=+1$ as a pseudo-vector, same as the EM background.
In addition, we are interested in the strengths of the induced electric and magnetic fields that
$E^{i}_{(1)}E_{i}^{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm sin}E_{i}^{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}~{}$
$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}-\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}+\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}+\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}$
which also works for $B^{i}_{(1)}B_{i}^{(1)}$.
We plot the strengths of electric and magnetic fields in different sections in FIG. 7, from which we can see that the strengths are mainly concentrated in $E^{2}_{\rm cos}$, $E^{2}_{\rm sin}$ and $B^{2}_{\rm sin}$, and for each panel there is a distinctive pattern of the distribution. It was mentioned in [31] that, when sources are present we can use the eigenfunctions of the angle-part equation to separate the master equation for $\varphi_{0}$ or $\varphi_{2}$ (their master equations without sources are separable) by expanding the source $T=\Sigma G(r)_{s}S_{lm}(\theta)e^{im\varphi}e^{-i\omega t}$. Remember that there are only two real degrees of freedom, i.e. once we get $\varphi_{0}$ or $\varphi_{2}$ then the other two complex quantities are decided. Therefore these featured patterns should come from the decomposition of the source, in our case it is the axion cloud but the expression of $T$ is so complicate to expand one by one.
Figure 7: The strengths of electric field and magnetic field $M^{2}E^{i}_{\rm cos}E^{\rm cos}_{i}$, $M^{2}E^{i}_{\rm sin}E^{\rm sin}_{i}$, $M^{2}E^{i}_{\rm cos}E^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm cos}_{i}$, $M^{2}B^{i}_{\rm sin}B^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm sin}_{i}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.2$.
VI Conclusions
In this work we studied the EM photon cloud sourced by an axion cloud through the axion-photon coupling around a Kerr black hole in the presence of a background EM field. In the first-order study of the coupling constant, we obtained an oscillating EM photon cloud with the same frequency as the axion cloud, which indicates that it exists even for very small couplings with no threshold. More intriguingly, this induced EM photon cloud is growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied.
In addition, we studied the induced electric and magnetic fields respectively, and found that their components exhibit very different symmetries in contrast with the EM background, which may indicate the existence of both an axion cloud and the axion-photon coupling. Furthermore, we analyzed the symmetries of most quantities in curved spacetime and confirmed the parity violation $P=-1$ of the $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ term. Besides, for near horizon region, we compared the distribution of the induced EM field with the boundary of photon region and found some overlap as expected.
Acknowledgements.
We thank Masahide Yamaguchi, Jin Sun, Yuhang Zhu and Cheng-Yong Zhang for beneficial discussions. This work is supported by IBS under the project code IBS-R018-D3.

### v4

@@
 In this work, we studied the EM photon cloud induced from an axion cloud around a Kerr black hole in the first order of the coupling constant classically.
-In the presence of a static EM background (like Wald extended solution) valid in realistic astrophysical environment, we found that an EM photon cloud emerges, oscillating at the same frequency as the axion cloud and growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied.
-Notably, the first-order study indicates that this photon cloud exists with no threshold for the coupling constant.
+In the presence of a static EM background (such as the extended Wald solution motivated by astrophysical environments), we found that an EM photon cloud emerges, oscillating at the same frequency as the axion cloud and growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied.
 The evolution of the EM photon cloud with time and azimuthal angle is obtained analytically while the cross-sectional distribution is solved numerically.
-Intriguingly, the induced EM field exhibits significantly different symmetries in contrast to the background EM field, which may serve as an indication of the existence of both an axion cloud and axion-photon coupling.
-Furthermore, in the near horizon region, the induced EM field has some overlap with the photon region as expected.
+The induced EM field exhibits symmetries that are markedly different from those of the background EM field.
+Consequently, the induced photon cloud forms an unstable bound configuration that emits EM waves to spatial infinity while being replenished by the axion cloud, providing a potential observational signature of both the presence of an axion cloud and axion-photon coupling.
 I Introduction
@@
 In (3+1)-dimensional case, a type called string-model independent axion [7, 8] is dual to the field strength of the Kalb-Ramond, spin-one, antisymmetric tensor field, and plays the role of a totally antisymmetric torsion in the geometry.
-Recently there are extensive experimental and observational efforts to find axionic imprints.
-Superradiant mechanisms can give us information about their existance as axions form dense clouds around spinning black holes [9].
-The superrradiance mechanism is a very well known method of extracting energy from a rotating black hole [10, 11, 12].
-It was shown by Penrose [13] that if a particle accretes into a Kerr black hole then energy can be extracted from the black hole.
-It was shown by Misner [14] that there is a constrain of the frequency of scatter waves off the black hole to be less than the rotational frequency of the black hole.
-Then if this constrained condition is satisfied it was shown by Teukolsky [15] that this energy extraction mechanism also works in the case of EM and gravitational waves, though the Kerr solution remains stable under massless scalar, EM and gravitational perturbations without reflection.
-It is well known that an isolated black hole cannot have an EM field unless it is endowed with a net electric charge.
-Therefore, for an isolated black hole to have some EM effects there should be a mechanism for charging it.
-If a black hole is in an astrophysical environment with a galactic magnetic field, EMs produced by external sources may charge it.
-In [16] an axisymmetric black hole which was placed in an originally uniform magnetic field and no significant field enhancement effects were found.
-However, if the black hole is rotating, the rotational effects of the black hole, produce electric fields near the black hole.
-Extraction of energy and charge from a black hole was studied in [17].
+Recently, there have been extensive experimental and observational efforts to search for axionic imprints.
+Superradiant mechanisms provide a promising avenue for probing their existence through gravitational waves [9, 10, 11], as axions can form dense clouds around spinning black holes.
+Although tidal perturbations from a nearby binary companion can destabilize the boson cloud and even terminate superradiance [12], the binary evolution can be significantly modulated by the backreaction with reduced termination rate [13].
+The superrradiance mechanism is a well established process for extracting energy from a rotating black hole [15, 14].
+Penrose and Floyd [16] first demonstrated that a particle entering the ergoregion of a rotating black hole can split such that one fragment falls in with negative energy, allowing the other to escape with excess energy, thereby extracting the black hole’s rotational energy.
+Subsequently, Misner [17] showed that waves with frequencies $\omega<m\Omega_{H}$ can also extract rotational energy through superradiant scattering, in which an incoming wave is amplified as it scatters off a rotating black hole.
+Later, Press and Teukolsky [18] introduced the concept of the black hole bomb, and Teukolsky further generalized the mechanism to EM and gravitational waves [19].
+Nevertheless, the Kerr solution remains stable against massless scalar, EM and gravitational perturbations without reflection.
+It is well known that an isolated black hole cannot sustain an intrinsic EM field unless it carries a net electric charge.
+Thus, for a black hole to exhibit EM effects, some charging mechanism is required.
+In realistic astrophysical environments, surrounding plasma and accretion flows in the presence of large-scale magnetic fields can give rise to effective charge separation or induce a nonvanishing net charge on the black hole.
+Possible origins of such EM configurations include charge separation induced by magnetospheric processes such as the Blandford–Znajek mechanism, effects associated with plasma inflow, disk–magnetosphere interactions, as well as pair-production–supported magnetospheres.
+In the present work, however, we do not model these mechanisms explicitly and instead adopt the extended Wald solution [20] as an effective, stationary background description.
+Neglecting backreaction and making use of the two Killing vectors of the Kerr spacetime, Wald constructed an exact test solution of Maxwell’s equations describing a homogeneous magnetic field aligned with the black hole spin axis, which could lead to the accretion of charges from the surrounding plasma in the accretion disk or the interstellar medium.
+As a result, the black hole would acquire a charge and be described by a Kerr-Newman spacetime.
+However, at equilibrium with the original magnetic field, this EM configuration is also a source-free test solution on the Kerr metric, so that we can consider it as the EM background solution in our work.
+Besides, extraction of energy and charge from a black hole was studied in [21].
 The superradiance effects and the multimessenger signals of an EM field were studied in [22, 11].
-It was shown in [20] that the presence of light scalars like axions can trigger superradiant instabilities around Kerr BHs and lead to the deposition of the black hole rotation energy in a cloud.
-Also it was shown in [21] that the presence of an axion field can lead to the reversal of rotation in Kerr black holes.
-It was also shown that instabilities are generated and a fraction of the cloud energy is transferred to EM blasts.
-The resulting large energy density of the cloud rotates around the black hole, and results to a monochromatic gravitational wave radiation that drain the cloud over parametrically longer times [22].
+Recently, new polarized images from the Event Horizon Telescope (EHT) collaboration revealed significant variations in brightness and polarization patterns across 2017, 2018, and 2021, highlighting the unexpectedly evolving magnetic field structures around the black hole M87* [23].
 Although the axion-photon coupling is expected to be very weak, it was suggested that the quantum stimulated emission of photons can be generated at large enough axion number [24, 25], and conjectured that blasts of light could be emitted from black hole systems [26, 27].
 Recently at classical level, it was shown that EM fields can be exponentially amplified in Kerr-axion cloud system, when the amplitude of the axion cloud multiplied by the coupling constant is larger than a critical value [28].
+Thereinto, a fraction of the cloud energy is transferred to the EM blasts, while the depleted axion cloud will be replenished via superradiance, leading to the spin-down of the black hole as it was found in [29].
+In [30], it was also found that the presence of an axion field can cause the reversal of rotation in Kerr black holes.
 Even in the presence of plasma, the EM instability is still controlled by the axionic coupling [31].
-In [27], the origin of the instability from the axion-photon coupling was analyzed in flat space through the Mathieu equation, and it was found that a background EM field can produce an axion profile around a Kerr black hole perturbatively.
-On the other side, from the European Horizon Telescope (EHT) observation of the polarimetric measurements of M87*, stringent constraints on the axion-photon coupling constant were given in the mass range of $\sim\left(10^{-21}-10^{-20}\right)$eV from the azimuthal distribution of electric vector position angle [28].
+Here the plasma effects are characterized by the plasma frequency $\omega_{p}$, which acts as an effective mass scale for EM waves and governs plasma screening.
+Interestingly, a new class of superradiant instabilities of axion bound states in neutron star magnetospheres has been identified [32], in which axion-photon mixing allows the extraction of rotational energy from the neutron star, since rotation and a dissipation channel are enough to trigger superradiance [15].
+In a flat-spacetime analysis with a uniform oscillating axion background, the instability appears through the Mathieu equation, originating from the homogeneous part of the equation [33].
+In contrast, the flat-spacetime analysis in a homogeneous EM background showed that the axion-photon coupling modifies the dispersion relation, and that a sufficiently strong electric field can trigger an instability with a growing mode.
+This latter instability arises from the inhomogeneous (sourced) part of the EM field equation.
+In fact, in the presence of external EM fields, photons can mix with low-mass bosons if they couple via a two-photon vertex, analogous to neutrino-flavor-mixing effects [34].
+In addition, both axion-photon mixing around charged black holes and dark photon-photon mixing around neutral black holes were studied in the presence of plasma [35], in which the flat spacetime analysis in the presence of a magnetic field gave the mixing matrixes with non-diagonal terms.
+Moreover, superradiant instabilities by photons confined around black holes in the presence of astrophysical plasmas were investigated under the plasma-photon interaction, where quasi-bound EM states were found to be superradiantly unstable for rotating black holes though it probably be quenched due to plasma blowout in the nonlinear evolution [36, 37, 38].
+From the observational side, with the polarimetric measurements of M87* by the EHT collaboration stringent constraints on the axion-photon coupling constant in the mass range of $\sim\left(10^{-21}-10^{-20}\right)$eV were given, based on the azimuthal distribution of the electric vector position angle [39].
 Furthermore, the deviation of photon geodesics due to oscillating metric perturbations induced from the superradiant clouds was predicted in [40].
@@
 It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part.
-It is surprising that even with very small couplings, no threshold is needed, there exists an oscillating photon cloud at the same frequency as the axion cloud and it will grow exponentially when the axion cloud is growing.
-Whereas in the previous studies, without a background EM field, a threshold for the amplitude of the axion cloud multiplied by the coupling constant and an initial EM pulse are both required, otherwise there is no instability and the initial EM fluctuation will decay exponentially.
-In contrast, for the dark photon-photon coupling case, a dark photon cloud can transfer its energy to form a rotating EM field directly [18].
+We found that no threshold on the coupling constant or on the amplitude of the axion cloud is required for the emergence of this oscillating and growing photon cloud.
+The instability is inherited from the axion cloud in the presence of an EM background.
+This is in contrast to the case without a background EM field, where a threshold for the amplitude of the axion cloud multiplied by the coupling constant and an initial EM pulse are both required; otherwise, no instability develops and the initial EM fluctuation decays exponentially.
+While for the dark photon-photon coupling case, a dark photon cloud can transfer its energy to form a rotating EM field directly without any external EM fields [22], since no background EM fields are necessary for the mixing to occur [35].
 The work is organized as follows.
@@
 Moreover, in Sec.
-IV we study the induced electric and magnetic fields respectively, and in Sec.
-V we analyze the symmetries of most quantities and confirm the parity violation.
+IV we study the induced electric and magnetic fields respectively.
 Finally in Sec.
-VI we conclude.
+V we conclude.
 II First-order Analysis of the Axion-Photon Coupling
@@
 $S=\int d^{4}x\sqrt{-g}\left[\frac{R}{2\kappa}+\mathcal{L}_{m}\right]~,$ (1)
-$\mathcal{L}_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\mu_{1}^{2}\phi^{2}-\frac{k_{a}}{2}\phi F_{\mu\nu}{}^{*}F^{\mu\nu}~{},$ (2)
-[…]
+$\mathcal{L}_{m}=-\frac{1}{4}F_{\mu\nu}F^{\mu\nu}-\frac{1}{2}\nabla_{\mu}\phi\nabla^{\mu}\phi-\frac{1}{2}\mu^{2}\phi^{2}-\frac{k_{a}}{2}\phi F_{\mu\nu}{}^{*}F^{\mu\nu}~,$ (2)
+where the axion field $\phi$ is a real pseudo scalar field with mass $m_{a}\equiv\mu\hbar$, $k_{a}$ is the coupling constant, $F_{\mu\nu}\equiv\nabla_{\mu}A_{\nu}-\nabla_{\nu}A_{\mu}$ is the electromagnetic field strength tensor and $^{*}F^{\mu\nu}\equiv\frac{1}{2}\epsilon^{\mu\nu\rho\sigma}F_{\rho\sigma}$ is the dual field strength tensor.
+Here $\epsilon^{\mu\nu\rho\sigma}\equiv\frac{1}{\sqrt{-g}}E^{\mu\nu\rho\sigma}$ and $E^{\mu\nu\rho\sigma}$ denotes the totally antisymmetric Levi-Civita symbol with $E^{0123}=+1$.
+We use geometrized units $c=G=1$ unless otherwise specified.
 By variation of the above action, we obtain the basic field equations
-$\nabla^{\mu}\nabla_{\mu}\phi-\mu_{1}^{2}\phi-\frac{k_{a}}{2}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~{},$ (3)
+$\nabla^{\mu}\nabla_{\mu}\phi-\mu^{2}\phi-\frac{k_{a}}{2}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~,$ (3)
 $\nabla_{\nu}F^{\mu\nu}+2k_{a}\nabla_{\nu}\phi{}^{*}F^{\mu\nu}=0~,$ (4)
@@
 In our study, the Kerr geometry is fixed as the background, since the backreactions of both axion cloud and the EMs on the metric are in the second-order of the fluctuations.
-We expand the scalar and EM field equations to the first-order of the coupling constant $k_{a}$
-$\nabla^{\mu}\nabla_{\mu}\phi_{(0)}-\mu_{1}^{2}\phi_{(0)}+k_{a}\left(\nabla^{\mu}\nabla_{\mu}\phi_{(1)}-\mu_{1}^{2}\phi_{(1)}-\frac{1}{2}F_{\mu\nu}^{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~{},$ (7)
+To clarify the hierarchy, we introduce small quantities $\eta$ and $e$ to represent the magnitudes of the axion cloud and the EM field respectively as $F_{\mu\nu}\to eF_{\mu\nu}$ and $\phi\to\eta\phi$, and the energy-momentum tensor becomes
+$T_{\mu\nu}$ $\to$ $-\frac{1}{2}g_{\mu\nu}\left(\frac{1}{2}e^{2}F_{\alpha\beta}F^{\alpha\beta}+\eta^{2}\nabla_{\sigma}\phi\nabla^{\sigma}\phi+\mu^{2}\eta^{2}\phi^{2}+k_{a}e^{2}\eta\phi F_{\alpha\beta}{}^{*}F^{\alpha\beta}\right)$~ (7)
+$+e^{2}F^{\sigma}{}_{\mu}F_{\sigma\nu}+\eta^{2}\nabla_{\mu}\phi\nabla_{\nu}\phi+4k_{a}e^{2}\eta\phi F^{\sigma}{}_{\nu}{}^{*}F_{\sigma\mu}~,$
+showing that the magnitudes of the backreaction on the metric are of the orders $\sim e^{2},~\eta^{2},~\mu^{2}\eta^{2},~k_{a}\eta e^{2}$. The other field equations become
+$\nabla^{\mu}\nabla_{\mu}\phi-\mu^{2}\phi-\frac{k_{a}e^{2}}{2\eta}F_{\mu\nu}{}^{*}F^{\mu\nu}=0~,$ (8)
+$\nabla_{\nu}F^{\mu\nu}+2k_{a}\eta\nabla_{\nu}\phi{}^{*}F^{\mu\nu}=0~,$ (9)
+which yields the conditions $\mu^{2}\eta\gg k_{a}e^{2}$ and $k_{a}\eta\ll 1$ when we consider the axion cloud and the EM field as background fields, with the effects of their interaction appearing only at higher orders.
+Using the typical field amplitude in the boson cloud $\langle\phi^{2}\rangle\sim 8\tilde{a}\alpha^{5}M_{\rm Pl}^{2}$ [41], and the limit for the largest magnetic field $B\sim 4\times 10^{18}\left(M/M_{\odot}\right)^{-1/2}{\rm Gauss}$ that can be supported by an accretion disk around astrophysical BHs [42], we obtain the first condition
+$k_{a}\ll\frac{160.705\sqrt{\tilde{a}}\alpha^{9/2}}{M_{\rm BH}/M_{\odot}}~{\rm GeV}^{-1}$~ (11)
+that the background solution of the axion cloud without EM field is valid, and the relevant axion mass is around
+$m_{a}\sim\frac{\alpha}{M_{\rm BH}/M_{\odot}}1.3\times 10^{10}~{\rm eV}~.$ (12)
+Here we adopt natural units, setting $\hbar=c=1$ and $4\pi\epsilon_{0}=1$, in order to facilitate comparison with other constraints on the coupling constant.
+For instance, in the case of the supermassive black hole M87*, taking $\tilde{a}=0.8$ and $\alpha=0.2$, we have $k_{a}\ll 1.58233\times 10^{-11}~{\rm GeV}^{-1}$.
+In contrast, for a primordial black hole with $M/M_{\odot}=0.2$, $\tilde{a}=0.5$ and $\alpha=0.2$, the condition becomes much less restrictive, $k_{a}\ll 0.406555~{\rm GeV}^{-1}$.
+We will verify the second validity condition of the perturbative framework for the EM field equation, after we obtain the amplitude of the first-order EM field.
+We now return to the original field equations and expand them to first order in the coupling constant $k_{a}$
+$\nabla^{\mu}\nabla_{\mu}\phi_{(0)}-\mu^{2}\phi_{(0)}+k_{a}\left(\nabla^{\mu}\nabla_{\mu}\phi_{(1)}-\mu^{2}\phi_{(1)}-\frac{1}{2}F_{\mu\nu}^{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~,$ (13)
 $\nabla_{\nu}F^{\mu\nu}_{(0)}+k_{a}\left(\nabla_{\nu}F^{\mu\nu}_{(1)}+2\nabla_{\nu}\phi_{(0)}{}^{*}F^{\mu\nu}_{(0)}\right)+\mathcal{O}\left(k_{a}^{2}\right)=0~,$ (14)
-which clearly shows that at zero-order the axion cloud and the EM field obey their own field equations respectively, and at first-order the existence of a background EM field can give a substructure on the axion cloud while the existence of both axion cloud and a background EM field can induce an extra EM field in order of $k_{a}$.
+which clearly shows that to zeroth order the axion cloud and the EM field obey their own field equations respectively, while at first order the background EM field gives a substructure on the axion cloud and the coexistence of the axion cloud and the background EM field generates an extra EM field of order $k_{a}$.
 The first-order backreaction on the axion cloud has been studied in [33].
@@
 which describes the equilibrium state of the black hole absorbing charges and discharging processes from the surroundings of astrophysical environments such as the accretion disk, plasma and the interstellar medium [20].
-It is worth noting that this extended Wald solution is still an exact solution in Kerr geometry.
-For the axion cloud part, the bound states in Kerr metric can be obtained analytically in approximation when the fine structure constant $\alpha_{1}\equiv GM\mu_{1}/c=M\mu_{1}\ll 1$ in separable form as
+It is worth noting that this extended Wald solution is still a source-free test solution on the Kerr geometry.
+For the axion cloud part, the bound states in Kerr metric can be obtained analytically in approximation when the fine structure constant $\alpha\equiv GM\mu/c=M\mu\ll 1$ in separable form as
 $\phi_{(0)}=e^{-i\omega t+im\varphi}S_{lm}(\theta)R_{lm}(r)$~ (20)
 with $\omega=\omega_{R}+i\omega_{I}$ that [43]
-$\omega_{R}$ $\simeq$ $\mu_{1}\left(1-\frac{\alpha_{1}^{2}}{2\left(n+l+1\right)^{2}}\right)~{},$ (15)
-$\omega_{I}$ $\simeq$ $\frac{1}{\gamma_{nlm}M}\left(\frac{am}{M}-2\mu_{1}r_{+}\right)\alpha_{1}^{4l+5}$ (16)
+$\omega_{R}$ $\simeq$ $\mu\left(1-\frac{\alpha^{2}}{2\left(n+l+1\right)^{2}}\right)~,$ (21)
+$\omega_{I}$ $\simeq$ $\frac{1}{\gamma_{nlm}M}\left(\frac{am}{M}-2\mu r_{+}\right)\alpha^{4l+5}$ (22)
 where $\gamma_{nlm}$ is a constant, $n$ is the principal quantum number, and $l$ corresponds to the orbital angular momentum quantum number.
-The condition for the superradiance to occur $\mu_{1}<\frac{am}{2Mr_{+}}=m\Omega_{H}$ is equivalent to $\alpha_{1}<\frac{am}{2r_{+}}=\frac{ma}{2M\left(1+\sqrt{1-a^{2}/M^{2}}\right)}$.
+The condition for the superradiance to occur $\mu<\frac{am}{2Mr_{+}}=m\Omega_{H}$ is equivalent to $\alpha<\frac{am}{2r_{+}}=\frac{ma}{2M\left(1+\sqrt{1-a^{2}/M^{2}}\right)}$.
 As we can see in eq.
@@
 It is well known that the timescale of the superradiant instability is much larger than the timescale of the oscillation, therefore we can ignore the growth of the amplitude and focus on the oscillating part.
-It is necessary to clarify the difference of the EM frequencies.
-Turning back to the original Klein-Gordon equation (3), an axion field evolving as $\phi\sim e^{-i\omega t+im\varphi}$ would lead to EM field as $F_{\mu\nu}\sim e^{-i\omega_{EM}t+im_{EM}\varphi}$ at half frequency of the axion field $\omega_{EM}=\omega/2$.
-While if we look at the EM field equation (4), we cannot get a direct relation between the axion frequency and the EM frequency.
-In quantum stimulated case [23], the half frequency was explained as the axion decay channel is into two photons resulting the photon energy at $E=m_{a}/2$.
-Also, the numerical evolution results [20, 9] gave EM oscillation frequency at half the axion mass frequency for supercritical couplings.
-In [27], the origin of the instability was analyzed in flat space through the Mathieu equation, which predicts instabilities for $\omega_{EM}=\mu_{1}/2,\mu_{1},3\mu_{1}/2,...$.
-However in [9] the frequency peaks at $\mu_{1},2\mu_{1},...$ were absent, and for high axionic couplings in dense plasmas $\omega_{p}>\mu_{1}/2$ the frequency centers at $\omega_{\rm EM}=\mu_{1}$ instead of the usual $\omega_{\rm EM}=\mu_{1}/2$, suggesting that the axion decay mechanism into two photons is forbidden and changes to process $\phi+\phi\to\gamma+\gamma$.
-Our understanding is that the half frequency appearing in classically numerical studies is because the main energy transfer from the axion cloud to EM field happens through the Klein-Gordon equation (3), so that even in curve spacetime we can still have a simple relation observed directly from the equation as we have analyzed previously.
+It is necessary to clarify the difference in the EM frequencies.
+Returning to the original Klein-Gordon equation (3), an axion field evolving as $\phi\sim e^{-i\omega t+im\varphi}$ would lead to the EM field as $F_{\mu\nu}\sim e^{-i\omega_{\rm EM}t+im_{\rm EM}\varphi}$ at half frequency of the axion field $\omega_{\rm EM}=\omega/2$.
+However, if we look at the EM field equation (4), a direct relation between the axion and the EM frequencies cannot be obtained.
+In the quantum stimulated case [24], the half frequency was explained by the axion decay channel $\phi\to 2\gamma$, where each photon carries energy $E=m_{a}/2$.
+Numerical simulations [28, 31] also showed that the EM oscillation frequency appears at half the axion mass frequency for supercritical couplings.
+In [33], the origin of the instability was analyzed in flat spacetime through the Mathieu equation, which predicts instabilities for $\omega_{\rm EM}=\mu/2,\mu,3\mu/2,$....
+Whereas in [31] the frequency peaks at $\mu,2\mu,$...
+were absent, and for high axionic couplings in dense plasmas with plasma frequency $\omega_{p}>\mu/2$, the EM spectrum instead centers at $\omega_{\rm EM}=\mu$ rather than at the usual $\omega_{\rm EM}=\mu/2$.
+This behavior suggests that the standard axion decay channel into two photons is kinematically forbidden and changes to process $\phi+\phi\to\gamma+\gamma$.
 III A Photon Cloud Induced from an axion cloud
-Next, we try to solve the first-order induced EM field $F^{\mu\nu}_{(1)}$ in the Newman-Penrose (NP) formalism, in which using the Kinnersley’s null tetrad [30]
+In this Section we will solve the first-order induced EM field $F^{\mu\nu}_{(1)}$ in the Newman-Penrose (NP) formalism, in which using the Kinnersley’s null tetrad [44]
 $l^{\mu}=\left(\frac{r^{2}+a^{2}}{\Delta_{r}},1,0,\frac{a}{\Delta_{r}}\right)~,~n^{\mu}=\frac{1}{2\Sigma}\left(r^{2}+a^{2},-\Delta_{r},0,a\right)~,~m^{\mu}=\frac{1}{\sqrt{2}\left(r+ia\cos{\theta}\right)}\left(ia\sin{\theta},0,1,\frac{i}{\sin{\theta}}\right)~,$ (23)
@@
 $\varphi_{0}=F_{\mu\nu}l^{\mu}m^{\nu}~,\quad\varphi_{1}=\frac{1}{2}F_{\mu\nu}\left(l^{\mu}n^{\nu}+\overline{m}^{\mu}m^{\nu}\right)~,\quad\varphi_{2}=F_{\mu\nu}\overline{m}^{\mu}n^{\nu}~,$ (24)
-where $\varphi_{0}$ is the ingoing radiation with spin weight $s=1$ and $\varphi_{2}$ represents outgoing radiation with spin weight $s=-1$.
+where $\varphi_{0}$ is the ingoing radiation with spin weight $s=1$ and $\varphi_{2}$ represents outgoing radiation with spin weight s=1s=-1.
 Note that there are only two real degree of freedoms though the three components are complex.
@@
 where $\psi=\varphi_{1}/\rho$, $J_{l}=J_{\mu}l^{\mu}$ and $J_{m}=J_{\mu}m^{\mu}$. The other two decoupled equations for $\varphi_{0}$ and $\varphi_{2}$ with sources can be found in [45].
-In fact, in the scenario we are considering, the axion cloud is the source for the induced EM field that
+In fact, in our framework, the axion cloud is the source for the induced EM field
 $\nabla_{\nu}F^{\mu\nu}_{(1)}=-2\nabla_{\nu}\phi_{(0)}{}^{*}F^{\mu\nu}_{(0)}\equiv J^{\mu}_{(1)}~.$ (32)
 We can take a good approximation to the bound states of axion cloud around a rotating black hole [14]
-$\phi_{(0)}=\alpha_{1}^{2}\phi_{0}\frac{r}{M}e^{-\frac{r}{2M}\alpha_{1}^{2}}\cos{\left(\varphi-\omega_{R}t\right)}\sin{\theta}~{},$ (27)
-which is valid for $\alpha\lesssim 0.2$ even at large black hole spin, and $\omega_{R}\simeq\mu_{1}$ is the oscillating frequency.
-It can be separated into two parts
+$\phi_{(0)}=\alpha^{2}\phi_{0}\frac{r}{M}e^{-\frac{r}{2M}\alpha^{2}}\cos{\left(\varphi-\omega_{R}t\right)}\sin{\theta}~,$ (33)
+which is valid for $\alpha\lesssim 0.2$ even at large black hole spin, and $\omega_{R}\simeq\mu$ is the oscillating frequency. It can be separated into two parts
 $\phi_{1}=e^{-i\omega t+i\varphi}R(r)S(\theta)~,\quad\phi_{2}=e^{i\omega t-i\varphi}R(r)S(\theta)~,$ (34)
-with $R(r)=\alpha_{1}^{2}\phi_{0}\frac{r}{2M}e^{-\frac{r}{2M}\alpha_{1}^{2}}$ and $S(\theta)=\sin{\theta}$.
+with $R(r)=\alpha^{2}\phi_{0}\frac{r}{2M}e^{-\frac{r}{2M}\alpha^{2}}$ and $S(\theta)=\sin{\theta}$.
 Using $\phi_{1}$ as the source and the NP scalars as partly separable variables
@@
 We can also do that for $\phi_{2}$.
-We replace $\omega=\mu_{1}=\alpha_{1}/M$, $a_{0}\equiv a/M$, $r_{0}\equiv r/M$, and define new dimensionless quantities $f_{01M}\equiv Mf_{01}\left(r,\theta\right)$, $f_{11M}\equiv Mf_{11}\left(r,\theta\right)$, $f_{\rm 21M}\equiv Mf_{21}\left(r,\theta\right)$.
-After absorbing some coefficients from the source part into the coupling constant $k_{a}$, we now use a new dimensionless coupling constant $\varepsilon\equiv\alpha_{1}^{2}B_{0}M\phi_{0}k_{a}$ as the small quantity for expansion, then the new field equations are dimensionless
-${\rm eq}f_{\rm 01M}:\left\{\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 01M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (30)
-${\rm eq}f_{\rm 11M}:\left\{\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 11M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (31)
-${\rm eq}f_{\rm 21M}:\left\{\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 21M}\left(r_{0},\theta\right),a_{0},\alpha_{1}\right\}~{},$ (32)
-and there are only two dimensionless parameters $a_{0}$ and $\alpha_{1}$.
+Then we replace $\omega=\mu=\alpha/M$, $a_{0}\equiv a/M$, $r_{0}\equiv r/M$, and define new dimensionless quantities $f_{01M}\equiv Mf_{01}\left(r,\theta\right)$, $f_{11M}\equiv Mf_{11}\left(r,\theta\right)$, $f_{\rm 21M}\equiv Mf_{21}\left(r,\theta\right)$.
+After absorbing some coefficients from the source part into the coupling constant $k_{a}$, we now use a new dimensionless coupling constant $\varepsilon\equiv\alpha^{2}B_{0}M\phi_{0}k_{a}$ as the small quantity for expansion, then the new field equations are dimensionless
+${\rm eq}f_{\rm 01M}:\left\{\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 01M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 01M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (36)
+${\rm eq}f_{\rm 11M}:\left\{\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 11M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 11M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (37)
+${\rm eq}f_{\rm 21M}:\left\{\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial r_{0}},\frac{\partial^{2}f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta^{2}},\frac{\partial f_{\rm 21M}\left(r_{0},\theta\right)}{\partial\theta},f_{\rm 21M}\left(r_{0},\theta\right),a_{0},\alpha\right\}~,$ (38)
+and there are only two dimensionless parameters $a_{0}$ and $\alpha$.
 Different from the quasi-normal modes and bound states, here the frequency of the induced EM field is already fixed by the frequency of the source (axion cloud), it is not an eigenvalue problem and no boundary condition should be imposed.
 Actually we can obtain the asymptotic behaviors of the functions at the boundaries by expanding the equations (we choose decaying solutions at spatial infinity), and then solve the equations numerically.
-With expansions and the regular condition at poles $\theta=0,\pi$, the asymptotic behaviors at the boundaries $r_{0}=r_{\rm h0}\equiv 1+\sqrt{1-a_{0}^{2}}$, $r_{0}\to\infty$, $\theta=0,\pi$ are as follows
-$f_{\rm 01M}\left(r_{\rm h0},\theta\right)=\frac{2\sqrt{2}a_{0}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{\theta}\sin{2\theta}}{\left(2\alpha_{1}r_{h0}-a_{0}-2i\sqrt{1-a_{0}^{2}}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (33)
-$f_{\rm 11M}\left(r_{\rm h0},\theta\right)=\frac{2i\sqrt{1-a_{0}^{2}}r_{h0}^{3}e^{-\alpha_{1}^{2}r_{\rm h0}/2}\sin{2\theta}}{\left(a_{0}-2\alpha_{1}r_{h0}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~{},$ (34)
-$f_{\rm 21M}\left(r_{\rm h0},\theta\right)=0~{},$ (35)
-$f_{\rm 01M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (36)
-$f_{\rm 11M}\left(r_{0}\to\infty,\theta\right)=-\frac{(\alpha_{1}+3i)\sin{2\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{4\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (37)
-$f_{\rm 21M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha_{1}^{2}r_{0}/2}}{2\sqrt{2}\alpha_{1}}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha_{1}^{2}r_{0}/2}\right)~{},$ (38)
-$f_{\rm 01M}\left(r_{0},0\right)=0,\quad f_{\rm 11M}\left(r_{0},0\right)=0,\quad f_{\rm 21M}\left(r_{0},0\right)=0,$ (39)
-$f_{\rm 01M}\left(r_{0},\pi\right)=0,\quad f_{\rm 11M}\left(r_{0},\pi\right)=0,\quad f_{\rm 21M}\left(r_{0},\pi\right)=0~{},$ (40)
-where the outgoing part $f_{\rm 21M}\left(r_{0},\theta\right)$ just vanishes at the horizon in accord with the fact that there is no outgoing radiation at the event horizon.
-It should also result from the fact that $\varphi_{2}$ is with negative spin weight which can not retrograde against the rotation at event horizon.
-In addition, it is clear that the asymptotic behaviors for $f_{\rm 01M}$ and $f_{\rm 11M}$ at event horizon $r_{\rm h0}$ are only valid when $\alpha_{1}\neq\frac{a_{0}}{2r_{\rm h0}}$, which is exactly the critical condition for whether superradiance occurs.
-In this critical case with $\alpha_{1}=\frac{a_{0}}{2r_{\rm h0}}$, the imaginary part $\omega_{I}$ disappears in the original field equation for the axion field, so that the whole frequency is not complex and the original boundary condition at event horizon is not valid, in which case another regular condition at origin should be applied.
-Here we do not go further on this case.
+The asymptotic behaviors of $f_{01M}$, $f_{11M}$ and $f_{\rm 21M}$ at the boundaries can be found in Appendix A.
 Similarly, for the NP scalars induced from $\phi_{2}$
@@
 These symmetries from analysis are consistent with the numerical results, thus here we only exhibit the distribution for those from $\phi_{1}$ in FIG.
-1.
-Figure 1: The real and imaginal parts of $f_{01M}\left(r_{0},\theta\right)$, $f_{11M}\left(r_{0},\theta\right)$, $f_{21M}\left(r_{0},\theta\right)$ induced from $\phi_{1}$ are plotted with $a_{0}=0.5$ and $\alpha_{1}=0.1$ in the scale of $200M$, where the superradiant condition is satisfied $\alpha_{1}<\frac{a_{0}}{2r_{\rm h0}}$.
+1, where the axes $x=r_{0}\sin{\theta}$ and $y=r_{0}\cos{\theta}$ are defined to describe the cross section.
+Figure 1: The real and imaginal parts of $f_{01M}\left(r_{0},\theta\right)$, $f_{11M}\left(r_{0},\theta\right)$, $f_{21M}\left(r_{0},\theta\right)$ induced from $\phi_{1}$ are plotted with $a_{0}=0.5$ and $\alpha=0.1$ in the scale of $200M$, where the superradiant condition is satisfied $\alpha<\frac{a_{0}}{2r_{\rm h0}}$.
 From FIG.
@@
 Usually the residential $\varphi_{1}$ part is neglected, since its equation is not separable even without any source and it is not related to ingoing or outgoing part.
-But here, it is shown that the imaginary part of $f_{11M}$ gets much larger amplitude than the other parts, indicating that the residential part is the primary constituent of the EM photon cloud we are studying.
-In fact, both $\varphi_{0}$ and $\varphi_{2}$ contain ingoing and outgoing parts, people assign ingoing/outgoing radiation to $\varphi_{0}$/$\varphi_{2}$ due to which one is dominant at spatial infinity.
-Subsequently, we can construct the EM tensor induced from $\phi_{1}$ using (19)
-$F_{\mu\nu}^{1}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]-4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{1}\left(r,\theta\right)\right]~{},$ (45)
-$P_{\mu\nu}^{1}\left(r,\theta\right)\equiv f_{01}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{11}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{11}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{21}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{}.$ (46)
-Similarly, for the EM tensor induced from $\phi_{2}$ we have
-$F_{\mu\nu}^{2}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]+4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}^{2}\left(r,\theta\right)\right]~{},$ (47)
-$P_{\mu\nu}^{2}\left(r,\theta\right)\equiv f_{02}\left(r,\theta\right)\overline{m}_{[\mu}n_{\nu]}+f_{12}\left(r,\theta\right)n_{[\mu}l_{\nu]}+f_{12}\left(r,\theta\right)m_{[\mu}\overline{m}_{\nu]}+f_{22}\left(r,\theta\right)l_{[\mu}m_{\nu]}~{},$ (48)
-then the whole EM field induced from the axion cloud $\phi_{(0)}=\phi_{1}+\phi_{2}$ evolves as
-$F_{\mu\nu}^{(1)}=F_{\mu\nu}^{1}+F_{\mu\nu}^{2}=\cos{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}\left(r,\theta\right)~{},$ (49)
+But here, it is shown that the imaginary part of $f_{11M}$ gets much larger amplitude than the other parts, indicating that the residential part is the primary constituent of the EM photon cloud we are concerned with.
+In fact, both $\varphi_{0}$ and $\varphi_{2}$ contain ingoing and outgoing parts, people assign ingoing/outgoing radiation to $\varphi_{0}$$\varphi_{2}$ due to which one is dominant at spatial infinity.
+Subsequently, we can construct the whole EM tensor induced from $\phi_{1}$ and $\phi_{2}$ as
+$F_{\mu\nu}^{(1)}=\cos{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}\left(r,\theta\right)~,$ (43)
+where
 $F_{\mu\nu}^{\rm cos}\left(r,\theta\right)$ $=$ $4~{\rm Re}\left[\left(f_{01}+f_{02}\right)\overline{m}_{[\mu}n_{\nu]}+\left(f_{11}+f_{12}\right)\left(n_{[\mu}l_{\nu]}+m_{[\mu}\overline{m}_{\nu]}\right)+\left(f_{21}+f_{22}\right)l_{[\mu}m_{\nu]}\right]~,$ (44)
@@
 This is one of the main results of our work.
-Moreover, it means that for an arbitrary current $J^{a}=e^{\pm i\omega t\mp im\varphi}S\left(r,\theta\right)$ as the source we will always get a photon cloud evolving as
-$F_{\mu\nu}=4\cos{\left(m\varphi-\omega t\right)}{\rm Re}\left[P_{\mu\nu}\left(r,\theta\right)\right]\pm 4\sin{\left(m\varphi-\omega t\right)}{\rm Im}\left[P_{\mu\nu}\left(r,\theta\right)\right]$.
-In fact, even for photons without any couplings, they can have an effective mass given by plasma and make superradiance to happen [32, 33].
+Moreover, it means that for an arbitrary current $J^{a}=\Sigma e^{\pm i\omega t\mp im\varphi}S\left(r,\theta\right)$ as the source we will always get a photon cloud evolving as (43).
+In fact, even photons without any couplings can acquire an effective mass in a plasma, allowing superradiance to occur for some primordial black holes [46, 41].
+In [31], a detailed study of the coupled axion-Maxwell system was performed on a black hole background, showing that, in the presence of plasmas, the EM instability remains controlled by the axion-photon coupling, even for relatively dense plasmas.
+In our case, the frequency of the induced photon cloud is twice the EM frequency in their work, making it less susceptible to plasma effects.
+Moreover, as shown in FIG.
+3, the induced EM scalar $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ is concentrated in regions away from both the equatorial plane and the polar axis, which renders it less affected by the accretion disk and relativistic jets.
+Importantly, the photon component of the cloud does not correspond to an on-shell propagating photon mode, but instead constitutes part of a mixed axion–photon propagation eigenstate undergoing superradiant amplification.
+Consequently, the existence of a photon component is not excluded by plasma screening.
+Furthermore, we emphasize that although the induced photon cloud is initially bound, it corresponds to unstable bound configurations that can subsequently escape to spatial infinity.
+Since photons experience no potential barrier at spatial infinity where the contribution from the axion cloud vanishes.
+From a particle interpretation, the photon component can be regarded as effectively free once it is released from the axion cloud into propagating EM modes.
+We now turn to the evolution of the induced EM photon cloud.
+The parity violation term $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ can be obtained as ($m=1$)
+$F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$~
+$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}~,$
+$F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$~
+$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}~,$
+which shows that except for the evolving parts with a double frequency $2\omega$, there are also residential parts that are time-independent.
+The $r,\theta$-dependent components $F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}$, $F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$, $F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$, $F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}$, $F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$ and $F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ are plotted in FIG.
+2-3.
+Their expressions can be found in Appendix B.
+Figure 2: The components $M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos},~M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin},~M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ of the parity violation term $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha=0.1,0.2$.
+Figure 3: The components $M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos},~M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin},~M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ of the contraction of the EM field strength tensor $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha=0.1,0.2$.
+We study the effects of the two parameters $a_{0},~\alpha$ and find that the spin parameter $a_{0}$ only has slight impacts on both the amplitude and the structure of the distribution for each components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$, while these quantities are sensitive to the fine structure constant $\alpha$.
+Therefore, in FIG.
+2-3 we plot the figures for the components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ with fixed $a_{0}=0.8$ and different values of $\alpha$.
+From the figures we can see that, for all the components, the increase of $\alpha$ leads to a significant decrease in amplitude and a slight concentration on the distribution, which indicates that for fixed black hole mass $M$ the smaller axion mass $\mu$ (larger reduced Compton wavelength) makes the interaction with photons more efficient and stronger.
+This is consistent with the idea that very light (or massless) particles may have a near-degeneracy with the photons, leading to a mixing phenomenon in which a coherent superposition of the two arises [34].
+Also, the slight concentration on the distribution should result from the concentration of the axion cloud with larger $\alpha$, where larger axion mass $\mu$ (smaller reduced Compton wavelength) makes the axion cloud concentrate.
+In general, the induced EM scalars $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ concentrate in the sine parts $M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ and $M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$.
+It is essential to verify the validity range of the perturbative framework with respect to the EM field, ensuring that $F_{\mu\nu}^{(0)}\gg\varepsilon F_{\mu\nu}^{(1)}$.
+Using $F_{\mu\nu}^{(0)}\sim B_{0}$, this requirement leads to
+$\alpha^{2}\phi_{0}k_{a}MF_{\mu\nu}^{(1)}\ll 1~.$ (48)
+For the case with $a_{0}=0.8$ and $\alpha=0.1$, the maximum value of $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ reaches approximately $8\times 10^{4}$, which gives $\phi_{0}k_{a}\ll 0.35$.
+While for the case with $a_{0}=0.8$ and $\alpha=0.2$, the constraint is significantly relaxed as $\phi_{0}k_{a}\ll 1.25$.
+Note that the quantity $\phi_{0}k_{a}$ is dimensionless in both geometrized and natural units.
+Remarkably, the resulting constraint is independent of the background field strength $B_{0}$.
 IV The induced electric and magnetic fields
@@
 $E^{a}:=n_{b}F^{ab}~,\quad B^{a}:=n_{b}{}^{*}F^{ab}~,$ (50)
-with respect to an Eulerian observer $n^{a}=\left(1/\alpha,-\beta^{i}/\alpha\right)$ whose worldline is orthogonal to the spacelike hypersurface.
-Therefore we have
+with respect to an Eulerian observer $n^{a}=\left(1/\alpha,-\beta^{i}/\alpha\right)$ whose worldline is orthogonal to the spacelike hypersurface. Therefore we have
 $E^{i}_{(1)}=\cos{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}\left(r,\theta\right)+\sin{\left(m\varphi-\omega t\right)}E^{i}_{\rm sin}\left(r,\theta\right)~,$ (51)
@@
 $B^{i}_{\rm cos}=-\frac{1}{2}\alpha\epsilon^{i0jk}F_{jk}^{\rm cos}~,\quad B^{i}_{\rm sin}=-\frac{1}{2}\alpha\epsilon^{i0jk}F_{jk}^{\rm sin}~.$ (54)
+The symmetries of these components are analyzed in Appendix C, consistent with the numerical results shown in FIG. 4-5.
 In FIG.
-2-4, we plot the components of induced electric and magnetic fields in cosine and sine sections respectively, in contrast with the background Wald case.
-In some panels with large scale like $100M$, we use the dotted circles to denote the radius $r_{\rm max}/M=2/\alpha_{1}^{2}$ of the maximum value of axion cloud, while in some panels with small scale like $5M$, we compare with the boundary of photon region (dotdashed curves) and event horizon (dashed circles).
+4-6, we plot the components of induced electric and magnetic fields in cosine and sine sections respectively, in contrast with the background Wald case.
+In some panels with large scale like $100M$, we use the dotted circles to denote the radius $r_{\rm max}/M=2/\alpha^{2}$ of the maximum value of axion cloud, while in some panels with small scale like $5M$, we compare with the boundary of photon region (dotdashed curves) and event horizon (dashed circles).
 It is a good case to compare the distribution of the EMs (in field perspective) with the photon region (spherical photon orbits in particle perspective).
-Figure 2: The components of the induced electric fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
-Figure 3: The components of the induced magnetic fields are plotted as cosine part and sine part respectively. In all the figures we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
-From the figures we can see that, the radial components $E^{r}_{\rm cos}$ and $E^{r}_{\rm sin}$ of the induced electric field both exhibit quadrupole moments with a parity violation ($P=-1$).
-In contrast, the radial component of electric field of the background Wald case is in dipole moments with a parity symmetry ($P=+1$).
-While the $\theta$ components $E^{\theta}_{\rm cos}$ and $E^{\theta}_{\rm sin}$ of the induced electric field both exhibit a parity symmetry ($P=+1$).
-In contrast, the $\theta$ component of electric field of the background Wald case is in quadrupole moments with a parity violation ($P=-1$).
-For the $\varphi$ components $E^{\varphi}_{\rm cos}$ and $E^{\varphi}_{\rm sin}$ of the induced electric field, they both show a parity violation symmetry ($P=-1$).
-Conversely, the symmetries for the components of magnetic fields are all opposite with those for electric fields.
+Figure 4: The dimensionless quantities $ME^{r}_{(1)}$, $M^{2}E^{\theta}_{(1)}$ and $M^{2}E^{\varphi}_{(1)}$ of the induced electric field components are plotted as cosine part and sine part respectively.
+In all the figures we have set $a_{0}=0.8$ and $\alpha=0.2$.
+Figure 5: The dimensionless quantities $MB^{r}_{(1)}$, $M^{2}B^{\theta}_{(1)}$ and $M^{2}B^{\varphi}_{(1)}$ of the induced magnetic field components are plotted as cosine part and sine part respectively.
+In all the figures we have set $a_{0}=0.8$ and $\alpha=0.2$.
+Figure 6: The dimensionless components $E^{r}_{\rm Wald}/B_{0}$, $ME^{\theta}_{\rm Wald}/B_{0}$, $B^{r}_{\rm Wald}/B_{0}$, $MB^{\theta}_{\rm Wald}/B_{0}$ of the electric and magnetic fields are plotted for the background Wald case, where we have set $a_{0}=0.8$.
+From the figures we can see that, the radial components $E^{r}_{\rm cos}$ and $E^{r}_{\rm sin}$ of the induced electric field both exhibit quadrupole structures with a sign reversal across the equatorial plane.
+In contrast, the radial component of the electric field in the background Wald solution is in dipole structure that is symmetric about the equatorial plane.
+While the $\theta$-components $E^{\theta}_{\rm cos}$ and $E^{\theta}_{\rm sin}$ of the induced electric field both exhibit even symmetry with respect to the equatorial plane, the $\theta$-component of the electric field in the background Wald solution shows a quadrupole pattern with odd symmetry.
+In addition, the $\varphi$-components $E^{\varphi}_{\rm cos}$ and $E^{\varphi}_{\rm sin}$ of the induced electric field possess odd symmetry about the equatorial plane.
+Conversely, the magnetic field components have opposite symmetries to those of the electric field components.
 Besides, we can observe some evident coincides between the distribution of the induced electric and magnetic fields and the boundary of photon region.
-Note that both electric and magnetic fields of the background Wald case do not contain $\varphi$ components, and they only distribute outside the ergoregion, while the induced EM field can live within the ergoregion.
+Note that both electric and magnetic fields of the background Wald case do not contain $\varphi$-components, and they only distribute outside the ergoregion, while the induced EM field can live within the ergoregion.
 This can be understood as “no static observer is allowed inside the ergoregion”.
-As we know, superradiance only happens for massive fields, where the mass term makes them gathering around the black hole and sustainably dissipate near the event horizon.
-The dissipation is a crucial ingredient for superradiance, in fact even for a rotating compact object without a horizon, superradiance can occur via other dissipation mechanism, like friction [10].
-While for the axion-photon coupling case, due to the antisymmetry of $\epsilon^{\mu\nu\rho\sigma}$ there is never an effective mass term for the photons, though sometimes people call it “tachyonic-like instability” [20].
-Also it has been mentioned that the time-dependent axion is a crucial ingredient to trigger this instability when $k_{a}\phi_{0}$ is larger than a critical value.
-In our first-order study with an EM background, even for small couplings, the exponential growing happens for the EM field coupled to the axion field, as we have discussed in Sec.
-II.
-From our perspective, the reason is that the stationary axion cloud ties the EM field onto the nearby of the black hole and serves as a source continuously, which makes the EM field dissipate sustainably from the axion cloud to the photon region and finally to the event horizon.
-As it is shown in the figures, the induced EM field does not concentrate around the maximum value of axion cloud but its main distribution is inside the dotted circles.
-In other words, the most dense places for the photon cloud do not relate to the largest oscillating position (same radius $r_{\rm max}$) of the axion cloud but inside it.
-Figure 4: The components of the electric and magnetic fields are plotted for the background Wald case, where we have set $a_{0}=0.8$ and $\alpha_{1}=0.2$.
-V Analysis of symmetries and parity violation
+Moreover, we are interested in the strengths of the induced electric and magnetic fields that
+$E^{i}_{(1)}E_{i}^{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm sin}E_{i}^{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}$~
+$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}-\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}+\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}+\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}$
+which also works for $B^{i}_{(1)}B_{i}^{(1)}$.
+Figure 7: The strengths of electric field and magnetic field $M^{2}E^{i}_{\rm cos}E^{\rm cos}_{i}$, $M^{2}E^{i}_{\rm sin}E^{\rm sin}_{i}$, $M^{2}E^{i}_{\rm cos}E^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm cos}_{i}$, $M^{2}B^{i}_{\rm sin}B^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm sin}_{i}$ are plotted with $a_{0}=0.8$ and $\alpha=0.2$.
+We plot the strengths of the electric and magnetic fields in different sections in FIG.
+7.
+It can be seen that the field strengths are mainly concentrated in $E^{2}_{\rm cos}$, $E^{2}_{\rm sin}$ and $B^{2}_{\rm sin}$, and that each panel exhibits a distinctive distribution pattern.
+As mentioned in [45], when sources are present, one can use the eigenfunctions of the angular equation to separate the master equation for $\varphi_{0}$ or $\varphi_{2}$ (their source-free master equations are separable) by expanding the source as $T=\Sigma G(r)_{s}S_{lm}(\theta)e^{im\varphi}e^{-i\omega t}$.
+Recall that there are only two real degrees of freedom; once $\varphi_{0}$ or $\varphi_{2}$ is determined, the other two complex quantities are fixed.
+Therefore, these featured patterns should come from the decomposition of the source.
+In our case, the source is the axion cloud, but the expression of $T$ is too complicate to expand term by term.
+As it is well known, superradiance occurs only for massive fields, where the mass term appears in the effective potential as an asymptotic constant at spatial infinity, providing a naturally reflective boundary that traps waves with lower oscillation frequencies.
+In addition, the mass term localizes the field around the black hole and allows it to sustainably dissipate near the event horizon.
+The dissipation is a crucial ingredient for superradiance.
+In fact, even for a rotating compact object without a horizon, superradiance can occur through other dissipation mechanisms, such as friction [15].
+In the case of axion-photon coupling, the antisymmetry of $\epsilon^{\mu\nu\rho\sigma}$ prevents the photons from acquiring an effective mass term for the photons; this is sometimes referred to as a “tachyonic-like instability” [28].
+It has also been noted that a time-dependent axion is crucial for triggering this instability when $k_{a}\phi_{0}$ exceeds a critical value.
+Furthermore, as mentioned earlier, in the presence of an external magnetic field, a flat spacetime analysis can give the mixing matrices for both the axion-photon coupling and the dark photon-photon coupling cases [35]
+$\mathcal{M}_{a\gamma}=\begin{pmatrix}-\omega_{p}^{2}/2\omega&B_{y}k_{a}\\ B_{y}k_{a}&-\mu_{a}^{2}/2\omega\end{pmatrix},\quad\mathcal{M}_{\gamma\gamma^{\prime}}=\frac{1}{2\omega}\begin{pmatrix}-\omega_{p}^{2}&\sin{\chi_{0}}\mu^{2}_{\gamma^{\prime}}\\ \sin{\chi_{0}}\mu^{2}_{\gamma^{\prime}}&-\mu_{\gamma^{\prime}}^{2}\end{pmatrix}$, (56)
+where the presence of a magnetic field enables the axion-photon coupling to effectively act as a “mass term” in the mixing, whereas the dark photon-photon mixing matrix does not require an external magnetic field.
+Moreover, the flat spacetime analysis shows that in a homogeneous EM background, the dispersion relation is modified by the axion-photon coupling [33]
+$\omega^{2}=\mathfrak{p}^{2}+\frac{1}{2}\mu^{2}\pm\frac{1}{2}\sqrt{\mu^{4}+16k_{a}^{2}E_{z}^{2}\mathfrak{p}^{2}}$~ (57)
+where a sufficiently large electric field leads to $\omega^{2}<0$, corresponding to a positive imaginary part of $\omega$ and thus a growing mode (instability).
+This is very similar to the superradiance case.
+Particularly the axions and photons share the same frequency, although the instability is driven by a stronger electric field, unlike our case.
+V Conclusions
+In this work we studied the EM photon cloud sourced by an axion cloud through the axion-photon coupling around a Kerr black hole in the presence of a background EM field.
+At first order in the coupling constant, we obtained an oscillating EM photon cloud with the same frequency as the axion cloud, indicating that it can exist even for very small couplings, with no threshold for either the coupling strength or the EM background field amplitude.
+More intriguingly, this induced EM photon cloud grows exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied within the perturbative framework.
+This instability is essentially inherited from the axion cloud, distinct from the instability that originates in the homogeneous part of the equation via the Mathieu equation, and also from that arising in the inhomogeneous (sourced) part of the EM field equation, where a sufficiently strong electric field is required.
+The growth and spatial distribution of the induced EM field are thus primarily governed by the axion cloud and the EM background field, though with various featured patterns.
+Both the EM background and the axion cloud serve as sources driving the growth of the induced EM photon cloud, which is mainly concentrated near the black hole while dissipating sustainably from the axion cloud to the photon region and finally toward the event horizon.
+As shown in the figures, the induced EM field does not peak at the location of the maximum axion density; rather, its main distribution lies within the dotted circles.
+In other words, the densest regions of the photon cloud do not coincide with the radius of maximum axion oscillation (same as $r_{\rm max}$) but are located inside it.
+In addition, we investigated the induced electric and magnetic fields respectively, and found that their components exhibit markedly different symmetries compared to those of the background EM field.
+As a result, the induced photon cloud forms an unstable bound configuration that emits EM waves to spatial infinity while being replenished by the axion cloud, providing a potential observational signature of both the presence of an axion cloud and axion-photon coupling.
+While a detailed comparison with black hole polarization images is beyond the scope of this work, the evolving EM structures induced by axion–photon coupling could, in principle, indirectly imprint subtle signatures in polarization observables—such as deviations from standard GRMHD polarization patterns or circular polarization—through their influence on plasma dynamics and radiative processes.
+Furthermore, we analyzed the symmetries of most relevant quantities, which are consistent with the numerically results shown in the figures.
+Besides, in the near-horizon region, we compared the distribution of the induced EM field with the boundary of photon region and found partial overlap, as expected.
+Acknowledgements.
+We thank Masahide Yamaguchi, Jin Sun, Yuhang Zhu, Minxi He and Cheng-Yong Zhang for beneficial discussions.
+We thank the editor and the referees for their constructive comments and suggestions.
+We also thank Enrico Cannizzaro for his comments, which improved our interpretations.
+Z.-Y.
+Tang is supported by IBS under the project code IBS-R018-D3.
+Appendix A Asymptotic behaviors of $f_{01M}$, $f_{11M}$ and $f_{\rm 21M}$
+With expansions and the regular condition at poles $\theta=0,\pi$, the asymptotic behaviors at the boundaries $r_{0}=r_{\rm h0}\equiv 1+\sqrt{1-a_{0}^{2}}$, $r_{0}\to\infty$, $\theta=0,\pi$ are as follows
+$f_{\rm 01M}\left(r_{\rm h0},\theta\right)=\frac{2\sqrt{2}a_{0}r_{h0}^{3}e^{-\alpha^{2}r_{\rm h0}/2}\sin{\theta}\sin{2\theta}}{\left(2\alpha r_{h0}-a_{0}-2i\sqrt{1-a_{0}^{2}}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~,$ (58)
+$f_{\rm 11M}\left(r_{\rm h0},\theta\right)=\frac{2i\sqrt{1-a_{0}^{2}}r_{h0}^{3}e^{-\alpha^{2}r_{\rm h0}/2}\sin{2\theta}}{\left(a_{0}-2\alpha r_{h0}\right)\left(r_{\rm h0}^{2}+a_{0}^{2}\cos^{2}{\theta}\right)^{2}}~,$ (59)
+$f_{\rm 21M}\left(r_{\rm h0},\theta\right)=0~,$ (60)
+$f_{\rm 01M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha^{2}r_{0}/2}}{\sqrt{2}\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (61)
+$f_{\rm 11M}\left(r_{0}\to\infty,\theta\right)=-\frac{(\alpha+3i)\sin{2\theta}e^{-\alpha^{2}r_{0}/2}}{4\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (62)
+$f_{\rm 21M}\left(r_{0}\to\infty,\theta\right)=-\frac{3i\sin^{2}{\theta}e^{-\alpha^{2}r_{0}/2}}{2\sqrt{2}\alpha}+\mathcal{O}\left(r_{0}^{-1}e^{-\alpha^{2}r_{0}/2}\right)~,$ (63)
+$f_{\rm 01M}\left(r_{0},0\right)=0,\quad f_{\rm 11M}\left(r_{0},0\right)=0,\quad f_{\rm 21M}\left(r_{0},0\right)=0,$ (64)
+$f_{\rm 01M}\left(r_{0},\pi\right)=0,\quad f_{\rm 11M}\left(r_{0},\pi\right)=0,\quad f_{\rm 21M}\left(r_{0},\pi\right)=0~,$ (65)
+where the outgoing part $f_{\rm 21M}\left(r_{0},\theta\right)$ just vanishes at the horizon in accord with the fact that there is no outgoing radiation at the event horizon.
+It should also result from the fact that $\varphi_{2}$ is with negative spin weight which can not retrograde against the rotation at event horizon.
+In addition, it is clear that the asymptotic behaviors for $f_{\rm 01M}$ and $f_{\rm 11M}$ at event horizon $r_{\rm h0}$ are only valid when $\alpha\neq\frac{a_{0}}{2r_{\rm h0}}$, which is exactly the critical condition for whether superradiance occurs.
+In this critical case with $\alpha=\frac{a_{0}}{2r_{\rm h0}}$, the imaginary part $\omega_{I}$ disappears in the original field equation for the axion field, so that the whole frequency is not complex and the original boundary condition at event horizon is not valid, in which case another regular condition at origin should be applied.
+Here we do not go further on this case.
+Appendix B Components of $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$
 It is well known that the term $F_{\mu\nu}{}^{*}F^{\mu\nu}$ violates parity and couples to the pseudo-scalar axion field, making the whole term $\phi F_{\mu\nu}{}^{*}F^{\mu\nu}$ parity invariant and leading to CP violation.
-Here in curved spacetime and classically, we obtain the same symmetry for the parity violation ($P=-1$). From (49), we have
-$F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}~{}$
-$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}~{},$
-where $m=1$ and
-$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}$ $=$ $8\left[({\rm Im}[f_{01}]+{\rm Im}[f_{02}])({\rm Re}[f_{21}]+{\rm Re}[f_{22}])+({\rm Re}[f_{01}]+{\rm Re}[f_{02}])({\rm Im}[f_{21}]+{\rm Im}[f_{22}])\right.$ (59)
+The components of the parity violation term $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$
+$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos}$ $=$ $8\left[({\rm Im}[f_{01}]+{\rm Im}[f_{02}])({\rm Re}[f_{21}]+{\rm Re}[f_{22}])+({\rm Re}[f_{01}]+{\rm Re}[f_{02}])({\rm Im}[f_{21}]+{\rm Im}[f_{22}])\right$.
+(66)
 $\left.-2({\rm Im}[f_{11}]+{\rm Im}[f_{12}])({\rm Re}[f_{11}]+{\rm Re}[f_{12}])\right]~,$
@@
 $+16({\rm Im}[f_{11}]-{\rm Im}[f_{12}])({\rm Re}[f_{11}]-{\rm Re}[f_{12}])~,$
-$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[-{\rm Im}[f_{01}]{\rm Im}[f_{21}]+{\rm Re}[f_{01}]{\rm Re}[f_{21}]+{\rm Im}[f_{02}]{\rm Im}[f_{22}]-{\rm Re}[f_{02}]{\rm Re}[f_{22}]\right.$ (61)
+$F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[-{\rm Im}[f_{01}]{\rm Im}[f_{21}]+{\rm Re}[f_{01}]{\rm Re}[f_{21}]+{\rm Im}[f_{02}]{\rm Im}[f_{22}]-{\rm Re}[f_{02}]{\rm Re}[f_{22}]\right$.
+(68)
 $\left.+{\rm Im}[f_{11}]^{2}-{\rm Re}[f_{11}]^{2}-{\rm Im}[f_{12}]^{2}+{\rm Re}[f_{12}]^{2}\right]~,$
 all exhibit a sign change for $\theta\to\pi-\theta$, in agreement with the results in FIG.
-5.
-Combined with no sign change for $\varphi\to\varphi+\pi$, we know that there is a parity violation $P=-1$ for $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$.
-In addition, it shows that except for the evolving parts with a double frequency $2\omega$, there are also residential parts that are time-independent.
-Figure 5: The components $M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}{}^{*}F^{\mu\nu}_{\rm sin}$ of the parity violation term $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
-Similarly, the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ evolves as
-$F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}~{}$
-$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}-\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}+\frac{1}{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}+\frac{1}{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}~{},$
-where $m=1$ and
+2.
+Combined with no sign change for $\varphi\to\varphi+\pi$, it shows a parity violation P=1P=-1 for $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$.
+While the components of the contraction of the EM field strength tensor $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$
 $F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos}$ $=$ $8\left[-({\rm Im}[f_{01}]+{\rm Im}[f_{02}])({\rm Im}[f_{21}]+{\rm Im}[f_{22}])+({\rm Re}[f_{01}]+{\rm Re}[f_{02}])({\rm Re}[f_{21}]+{\rm Re}[f_{22}])\right$.~ (69)
 $\left.+({\rm Im}[f_{11}]+{\rm Im}[f_{12}])^{2}-({\rm Re}[f_{11}]+{\rm Re}[f_{12}])^{2}\right]~,$
-$F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[({\rm Im}[f_{01}]-{\rm Im}[f_{02}])({\rm Im}[f_{21}]-{\rm Im}[f_{22}])-({\rm Re}[f_{01}]-{\rm Re}[f_{02}])({\rm Re}[f_{21}]-{\rm Re}[f_{22}])\right.$ (64)
+$F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$ $=$ $8\left[({\rm Im}[f_{01}]-{\rm Im}[f_{02}])({\rm Im}[f_{21}]-{\rm Im}[f_{22}])-({\rm Re}[f_{01}]-{\rm Re}[f_{02}])({\rm Re}[f_{21}]-{\rm Re}[f_{22}])\right$.
+(70)
 $\left.-({\rm Im}[f_{11}]-{\rm Im}[f_{12}])^{2}+({\rm Re}[f_{11}]-{\rm Re}[f_{12}])^{2}\right]~,$
@@
 all keep a parity symmetry $P=+1$ under $\theta\to\pi-\theta$, in agreement with the results in FIG.
-6.
+3.
 Also there is no sign change under $\varphi\to\varphi+\pi$, hence the whole $F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ keeps parity symmetry $P=+1$.
@@
 Here with the axion-photon coupling, it shows obviously that the null conditions are not satisfied anymore.
-Figure 6: The components $M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm cos},~{}M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin},~{}M^{2}F_{\mu\nu}^{\rm cos}F^{\mu\nu}_{\rm sin}$ of the contraction of the EM field strength tensor $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.1,0.2$.
-Moreover, we study the effects of the two parameters $a_{0},~{}\alpha_{1}$ and find that the spin parameter $a_{0}$ only has slight impacts on both the amplitude and the structure of the distribution for each components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$, while these quantities are sensitive to the fine structure constant $\alpha_{1}$.
-Therefore, in FIG.
-5-6 we plot the figures for the components of $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ with fixed $a_{0}=0.8$ and different values of $\alpha_{1}$.
-From the figures we can see that, for all the components, the increase of $\alpha_{1}$ leads to a significant decrease in amplitude and a slight concentration on the distribution, which indicates that for fixed black hole mass $M$ the smaller axion mass $\mu_{1}$ (larger reduced Compton wavelength) makes the interaction with photons more efficient and stronger.
-Also, the slight concentration on the distribution should result from the concentration of the axion cloud with larger $\alpha_{1}$, where larger axion mass $\mu_{1}$ (smaller reduced Compton wavelength) makes the axion cloud concentrate.
-In general, the induced EM scalars $M^{2}F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ and $M^{2}F_{\mu\nu}^{(1)}F^{\mu\nu}_{(1)}$ concentrate in the sine parts $M^{2}F_{\mu\nu}^{\rm sin}{}^{*}F^{\mu\nu}_{\rm sin}$ and $M^{2}F_{\mu\nu}^{\rm sin}F^{\mu\nu}_{\rm sin}$.
+Appendix C Symmetries of the electric and magnetic fields
 For the electric and magnetic fields respectively, applying the relations (40)-(42), we can obtain some symmetries for the components in cosine and sine parts
-$E^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{r}_{\rm cos}\left(r,\theta\right)~{},\quad B^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{r}_{\rm cos/sin}\left(r,\theta\right)~{},$ (66)
+$E^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{r}_{\rm cos/sin}\left(r,\theta\right)~,\quad B^{r}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{r}_{\rm cos/sin}\left(r,\theta\right)~,$ (72)
 $E^{\theta}_{\rm cos/sin}\left(r,\pi-\theta\right)=E^{\theta}_{\rm cos/sin}\left(r,\theta\right)~,\quad B^{\theta}_{\rm cos/sin}\left(r,\pi-\theta\right)=-B^{\theta}_{\rm cos/sin}\left(r,\theta\right)~,$ (73)
 $E^{\varphi}_{\rm cos/sin}\left(r,\pi-\theta\right)=-E^{\varphi}_{\rm cos/sin}\left(r,\theta\right)~,\quad B^{\varphi}_{\rm cos/sin}\left(r,\pi-\theta\right)=B^{\varphi}_{\rm cos/sin}\left(r,\theta\right)~,$ (74)
-which can be clearly seen in FIG. 2-3. We further have
+which can be clearly seen in FIG. 4-5. We further have
 $E^{r}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=E^{r}_{(1)}\left(t,r,\theta,\varphi\right)~,\quad B^{r}_{(1)}\left(t,r,\pi-\theta,\varphi+\pi\right)=-B^{r}_{(1)}\left(t,r,\theta,\varphi\right)~,$ (75)
@@
 These transformations are equivalent to $\overset{\rightharpoonup}{E}\rightarrow-\overset{\rightharpoonup}{E}$ and $\overset{\rightharpoonup}{B}\rightarrow\overset{\rightharpoonup}{B}$ under $\overset{\rightharpoonup}{r}\rightarrow-\overset{\rightharpoonup}{r}$.
-Therefore the induced electric field still keep a parity symmetry with $P=-1$ as a vector, and the induced magnetic field keeps the parity symmetry with $P=+1$ as a pseudo-vector, same as the EM background.
-In addition, we are interested in the strengths of the induced electric and magnetic fields that
-$E^{i}_{(1)}E_{i}^{(1)}$ $=$ $\cos^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm cos}+\sin^{2}{\left(m\varphi-\omega t\right)}E^{i}_{\rm sin}E_{i}^{\rm sin}+2\cos{\left(m\varphi-\omega t\right)}\sin{\left(m\varphi-\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}~{}$
-$=$ $\cos{\left(2m\varphi-2\omega t\right)}\left(\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}-\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}\right)+\sin{\left(2m\varphi-2\omega t\right)}E^{i}_{\rm cos}E_{i}^{\rm sin}+\frac{1}{2}E^{i}_{\rm cos}E_{i}^{\rm cos}+\frac{1}{2}E^{i}_{\rm sin}E_{i}^{\rm sin}$
-which also works for $B^{i}_{(1)}B_{i}^{(1)}$.
-We plot the strengths of electric and magnetic fields in different sections in FIG.
-7, from which we can see that the strengths are mainly concentrated in $E^{2}_{\rm cos}$, $E^{2}_{\rm sin}$ and $B^{2}_{\rm sin}$, and for each panel there is a distinctive pattern of the distribution.
-It was mentioned in [31] that, when sources are present we can use the eigenfunctions of the angle-part equation to separate the master equation for $\varphi_{0}$ or $\varphi_{2}$ (their master equations without sources are separable) by expanding the source $T=\Sigma G(r)_{s}S_{lm}(\theta)e^{im\varphi}e^{-i\omega t}$.
-Remember that there are only two real degrees of freedom, i.e.
-once we get $\varphi_{0}$ or $\varphi_{2}$ then the other two complex quantities are decided.
-Therefore these featured patterns should come from the decomposition of the source, in our case it is the axion cloud but the expression of $T$ is so complicate to expand one by one.
-Figure 7: The strengths of electric field and magnetic field $M^{2}E^{i}_{\rm cos}E^{\rm cos}_{i}$, $M^{2}E^{i}_{\rm sin}E^{\rm sin}_{i}$, $M^{2}E^{i}_{\rm cos}E^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm cos}_{i}$, $M^{2}B^{i}_{\rm sin}B^{\rm sin}_{i}$, $M^{2}B^{i}_{\rm cos}B^{\rm sin}_{i}$ are plotted with $a_{0}=0.8$ and $\alpha_{1}=0.2$.
-VI Conclusions
-In this work we studied the EM photon cloud sourced by an axion cloud through the axion-photon coupling around a Kerr black hole in the presence of a background EM field.
-In the first-order study of the coupling constant, we obtained an oscillating EM photon cloud with the same frequency as the axion cloud, which indicates that it exists even for very small couplings with no threshold.
-More intriguingly, this induced EM photon cloud is growing exponentially in accordance with the axion cloud when the superradiant condition for the axion field is satisfied.
-In addition, we studied the induced electric and magnetic fields respectively, and found that their components exhibit very different symmetries in contrast with the EM background, which may indicate the existence of both an axion cloud and the axion-photon coupling.
-Furthermore, we analyzed the symmetries of most quantities in curved spacetime and confirmed the parity violation $P=-1$ of the $F_{\mu\nu}^{(1)}{}^{*}F^{\mu\nu}_{(1)}$ term.
-Besides, for near horizon region, we compared the distribution of the induced EM field with the boundary of photon region and found some overlap as expected.
-Acknowledgements.
-We thank Masahide Yamaguchi, Jin Sun, Yuhang Zhu and Cheng-Yong Zhang for beneficial discussions. This work is supported by IBS under the project code IBS-R018-D3.
+Therefore the induced electric field still keep a parity symmetry with P=1P=-1 as a vector, and the induced magnetic field keeps the parity symmetry with $P=+1$ as a pseudo-vector, same as the EM background.